### v1

# Measuring deviations from spherical symmetry

Abstract
Most of the work on checking spherical symmetry assumptions on the distribution of $p$-dimensional random vector $Y$ has its focus on statistical tests for the null hypothesis of exact spherical symmetry. In this paper we take a different point of view and propose a measure for the deviation from spherical symmetry, which is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a random vector $Y_{s}$ with a spherical distribution. We develop estimators for the minimum distance with corresponding statistical guarantees (provided by asymptotic theory) and demonstrate the applicability of our approach by means of a simulation study and a real data example.
1 Introduction
Spherical contoured models include multivariate Gaussian distribution, multivariate $t$- distribution and multivariate stable distribution find their applications in earth science, vector cardiology (Mardia and Jupp,, 1999; Di Marzio et al.,, 2019), portfolio theory (Gupta et al.,, 2013), Mahalanobis distances (Luis Angel García-Escudero,, 2005). There has been a wide interest in testing for the spherical symmetry, or a wider class of elliptical symmetry, see for example Smith, (1977), Baringhaus, (1991), Fang et al., (1993), Koltchinskii and Li, (1998), Schott, (2002), Huffer and Park, (2007), Liang et al., (2008), Cassart et al., (2008), Henze et al., (2014), Albisetti et al., (2020), Babić et al., (2021), Banerjee and Ghosh, (2024) among others. A common feature of most of the cited references is that the authors use different approaches to construct tests for the null hypothesis of exact spherical symmetry. However, this assumption is rarely met because one believes that it is exactly (with mathematical equality) satisfied, but with the hope that the deviation from the null hypothesis is small such that (optimal) inference under the assumption of spherical symmetry is still reliable. Thus, strictly speaking, there are many situations where tests for exact sphericity are performed, although it is clear that this hypothesis is at most “approximately” satisfied.
In this paper, we take a different point of view on the problem of validating the assumption of spherical symmetry on the distribution of a random vector $Y$. Instead of testing for spherical symmetry, we propose a measure of deviation from spherical symmetry. This measure is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a vector $Y_{s}$ with a spherical distribution, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{p}$ (note that a random variable $Y_{s}$ has a spherical distribution if and only if $\|Y_{s}\|$ and $Y_{s}/\|Y_{s}\|$ are independent and $Y_{s}/\|Y_{s}\|$ is uniformly distributed on the $p$-dimensional sphere $\mathbb{S}_{p-1}$). In Section 2 we derive an explicit expression for this minimum distance, which is essentially the difference between the (squared) $L^{2}$-norm of the density of the vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ and the (squared) $L^{2}$-norm of the density of the random variable $\|Y_{s}\|$ multiplied with the surface area of $\mathbb{S}^{p-1}$. Consequently, the estimation problem of the minimum distance boils down to estimating the squared $L^{2}$-norm of the density of a distribution on the sphere and a distribution on $\mathbb{R}^{+}$. While the problem of estimating the integrated squared density on $\mathbb{R}^{p}$ has been extensively studied from various perspectives including optimality and adaptivity (see Hall and Marron,, 1987; Bickel and Ritov,, 1988; Laurent and Massart,, 2000; Giné and Nickl,, 2008, among many others), the corresponding problem on the sphere have not found much attention so far. We construct an estimator for the minimum distance from a sequence of independent identically distributed observations combining an estimator for the integrated squared density of the joint distribution of $Y/\|Y\|$ and $\|Y\|$ with an estimator for the integrated squared density of $\|Y\|$ on the non-negative line. We prove asymptotic normality of an appropriately centered and normalized version. Interestingly, from the theoretical perspective, there appear to be different scalings, depending on whether the distribution is spherical symmetric or not. In the latter case, the difference of the estimators converges at a rate $\sqrt{n}$ as in the case estimating the integral of the squared density (see, for example, Bickel and Ritov,, 1988). On the other hand, if the distribution of $Y$ is in fact spherical symmetric, the estimator converges at a rate $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, where $h\to 0$ and $\kappa\to\infty$ denote smoothing parameters required for the estimation of the integrated squared densities of $\|Y\|$ and $Y/\|Y\|$, respectively.
Based on these results, we develop several statistical applications. In Section 3.2 we derive an asymptotic confidence interval for the deviation of the distribution from sphericity. Moreover, we also construct tests for the hypothesis that the deviation from sphericity exceeds a given threshold. For example, if ${\cal M}^{2}$ denotes the measure of deviation we derive a consistent and asymptotic level $\alpha$-test for the hypotheses $H_{0}^{\rm eq}:{\cal M}^{2}>\Delta$ versus $H_{1}^{\rm eq}:{\cal M}^{2}\leq\Delta$, which allows to decide for approximate sphericity at a controlled type $I$ error. These procedures require the estimation of a complicated asymptotic variance, which we address by a jackknife procedure. Moreover, as an alternative, we also develop in Section 4 pivotal inference for the measure ${\cal M}^{2}$, which does not require any estimation of asymptotic variances. Our approach is based on the self-normalization principle (see Shao,, 2015), which cannot be used directly in the present context, but has to be adapted to the specific statistical problems under consideration. In particular, we have to prove the weak convergence of a sequential estimator of ${\cal M}^{2}$ to obtain a pivotal statistic which can be used for the construction of confidence intervals and for testing the hypotheses of a relevant deviation from sphericity. Finally, we demonstrate the applicability of our approach through a small simulation study and a real data example.
2 A measure of deviation and its estimation
Let $Y$ denote a $p$-dimensional random variable and define $U=\|Y\|$, $V=Y/U$, where $\|\cdot\|$ denotes the Euclidean norm. The density function of $Y$ can be written as
$f_{Y}(y)=f_{U,V}(u,v)=f_{U}(u)f_{V|U}(u,v),$ (2.1)
where $u=\|y\|$ and $v=y/\|y\|$. Throughout this paper, we will use $f$ for $f_{U,V}$. The distribution of $Y$ is called spherically distributed if and only if its density function can be represented in the form
$f_{Y}(y)=g(\|y\|),$ (2.2)
for some function $g$. Consequently, observing (2.1), it follows that $Y$ has a spherical symmetric distribution if and only $U=\|Y\|$ and $V=Y/\|Y\|$ are independent and $V=Y/\|Y\|$ is uniformly distributed on the sphere with (constant) density
$f_{0}(v)=\omega_{p-1}^{-1}I(\|v\|=1),$ (2.3)
where $\omega_{p-1}=2\pi^{p/2}/\Gamma(p/2)$ denotes the surface area of the sphere $\mathbb{S}^{p-1}$. Note that any spherical distribution of a $p$-dimensional random variable corresponds to a density of the form $f_{U}(u)f_{0}(v)$ for some density $f_{U}$ on $\mathbb{R}^{+}$, where $f_{0}$ is given by (2.3). Consequently, we define
$\mathcal{M}^{2}=\min_{h}\int_{\mathbb{R}^{+}}\int_{\mathbb{S}^{p-1}}(f(u,v)-h(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v),$ (2.4)
as a measure for the deviation from spherical symmetry, where $\omega_{p-1}(\mathrm{d}v)$ is the area element of $\mathbb{S}^{p-1}$, where the minimum is taken over the set of densities on $\mathbb{R}^{+}$. Note that $\omega_{p-1}=\int_{\mathbb{S}^{p-1}}\omega_{p-1}(\mathrm{d}v)$. Our first result provides an explicit solution of this optimization problem, and is proved in Section B.1.
Proposition 2.1.
The minimum in (2.4) is obtained for the marginal density of $(U,V)=(\|Y\|,Y/\|Y\|)$, that is $h^{*}(u):=f_{U}(u)=\int_{\mathbb{S}^{p-1}}f(u,v)\ \omega_{p-1}(\mathrm{d}v)$, and given by
$\mathcal{M}^{2}=\int_{\mathbb{R}^{+}}\int_{\mathbb{S}^{p-1}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)-\omega_{p-1}^{-1}\int_{\mathbb{R}^{+}}f^{2}_{U}(u)\mathrm{d}u.$ (2.5)
Let $Y_{1},\ldots,Y_{n}$ denote independent identically distributed $p$-dimensional random variables and define $U_{i}=\|Y_{i}\|$; $V_{i}=Y_{i}/\|Y_{i}\|$ ($i=1,\ldots,n)$. By Proposition 2.1 the problem of estimating the minimum distance ${\cal M}^{2}$ boils down to the estimation of the integrated squared densities of the random variables $(U_{i},V_{i})$ and $U_{i}$. The latter can be estimated by standard methods (see, for example Hall and Marron,, 1987; Bickel and Ritov,, 1988, among many others) and we use
$\hat{\cal M}_{n}^{(1)}:=\frac{1}{n(n-1)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)$ (2.6)
as estimate for the integrated squared density of $U$, where $h$ denotes a bandwidth. The estimation of the first integral in (2.5) is more intricate as it corresponds to the (squared) $L^{2}$-norm of a density on $\mathbb{R}^{+}\times\mathbb{S}^{p-1}$. For this purpose, we use classical results from density estimation on the sphere, such as Hall et al., (1987), and consider a rapidly varying function $L:\mathbb{R}^{p}\to\mathbb{R}$, which satisfies (i) $L$ is nonnegative and nondecreasing, (ii) for each $0<r<1$,
$L(rt)/L(t)\to 0,\ \text{as}\ t\to\infty.$ (2.7)
The estimator of the first integral in Proposition 2.1 is then defined by
$\hat{\cal M}_{n}^{(2)}:=\frac{1}{n(n-1)c_{1}(\kappa)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j})$ (2.8)
where $K$ is a kernel function and $h$ is the same bandwidth as used in the estimator $\hat{\cal M}_{n}^{(1)}$, $\kappa\to\infty$ is another bandwidth corresponding to the spherical part of the vector $(U,V)$ and the constant $c_{1}(\kappa)$ is given by
$c_{1}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta.$ (2.9)
Throughout this paper, we will consider the Fisher-von-Mises distribution function
$L(\kappa t)=\kappa^{p/2-1}\{(2\pi)^{p/2}\mathcal{I}_{p/2-1}(\kappa)\}^{-1}e^{\kappa t}$ (2.10)
for the kernel $L$, where $\mathcal{I}_{\nu}$ is the modified Bessel function of the first kind (see also 9.6.18 on Page 376 of Abramowitz and Stegun,, 1968) of order $\nu$, i.e.,
$\mathcal{I}_{\nu}(\kappa)=\frac{(\kappa/2)^{\nu}}{\Gamma(\nu+1/2)\Gamma(1/2)}\int_{-1}^{1}e^{\kappa t}(1-t^{2})^{\nu-1/2}\mathrm{d}t.$ (2.11)
Finally, we propose to estimate the minimum distance in (2.5) by the $U$-statistic
$\hat{\mathcal{M}}^{2}_{n}$ $=\hat{\mathcal{M}}^{(2)}_{n}-\omega_{p-1}^{-1}\hat{\mathcal{M}}^{(1)}_{n}=\frac{2}{n(n-1)}\sum_{i=1}^{n}\sum_{j=1}^{i-1}H_{n}(Y_{i},Y_{j}),$ (2.12)
where $\hat{\mathcal{M}}^{(1)}_{n}$ and $\hat{\mathcal{M}}^{(2)}_{n}$ are defined in (2.6) and (2.8), respectively, and
$H_{n}(Y_{i},Y_{j})=c_{1}^{-1}(\kappa)h^{-1}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j})-\omega_{p-1}^{-1}h^{-1}K\left(\frac{U_{i}-U_{j}}{h}\right)$ (2.13)
is a kernel of order $2$, where $U_{i}=\|Y_{i}\|$, $V_{i}=Y_{i}/\|Y_{i}\|$ ($i=1,\ldots,n)$. Note that the kernel depends on the sample size $n$ through the bandwidths $h$ and $\kappa$.
3 Asymptotic properties and first statistical applications
In this section, we establish a central limit theorem for the statistic $\hat{\mathcal{M}}_{n}^{2}$ as defined in (2.12). Interestingly, the convergence rate of $\hat{\mathcal{M}}^{2}_{n}$ to $\mathcal{M}^{2}$ depends on whether the distribution of $Y$ is spherically symmetric or not. We use these results to develop several statistical applications in Section 3.2.
3.1 Asymptotic theory
For the investigation of the bias of $\hat{\mathcal{M}}_{n}^{2}$, we recall a Taylor expansion for a twice differentiable function $f:\mathbb{R}^{+}\times\mathbb{S}^{p-1}\to\mathbb{R}$ with respect to the argument $v$ on the sphere $\mathbb{S}^{p-1}$. According to equation (3) in Marco Di Marzio and Taylor, (2014), we have
$f(u,y)=f(u,x)+\theta\xi^{\top}\mathcal{D}_{f}(u,x)+\frac{\theta^{2}}{2}\xi^{\top}\mathcal{D}^{2}_{f}(u,x)\xi+O(\theta^{3}).$ (3.1)
Here $D_{f}^{s}(u,x)$ is the matrix of the $s$th-order derivative of the function $f(u,v)$ with respect to $v$ at the point $x$ and for $x,y\in\mathbb{S}^{p-1}$, the vector $\xi$ and the angle $\theta\in(0,\pi)$ are defined by the tangent–normal decomposition
$y=x\cos(\theta)+\xi\sin(\theta).$ (3.2)
Note that $\theta\in(0,\pi)$ is the angle between $x$ and $y$, and $\xi\in\mathbb{S}^{p-1}$ is a vector orthogonal to $x$, i.e., $\xi\in\Omega_{x}=\{\xi\in\mathbb{S}^{p-1}:\xi\perp x\}$. The intuition for (3.1) comes from the fact that $y-x\approx\theta\xi$ when $\theta$ approximates 0, which follows from a Taylor expansion of the function $\cos\theta$ and $\sin\theta$ at the point $\theta=0$. Moreover, the transformation of the area element of $\mathbb{S}^{p-1}$ corresponding to the tangent–normal decomposition (3.2) is given by
$\omega_{p-1}(\mathrm{d}y)=(\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi),$ (3.3)
see equation (2) in Marco Di Marzio and Taylor, (2014) or equation (1.5) in Hall et al., (1987). On the other hand, for a fixed $v\in\mathbb{S}^{p-1}$, we have the Taylor expansion for $f$
$f(u_{0}+h,v)=f(u_{0},v)+h\left.\frac{\partial f(u,v)}{\partial u}\right|_{u=u_{0}}+\frac{h^{2}}{2}\left.\frac{\partial^{2}f(u,v)}{\partial u^{2}}\right|_{u=u_{0}}+O(h^{3}),$ (3.4)
where $\frac{\partial^{s}f(u,v)}{\partial u^{s}}$ denotes the partial derivative of order $s$ with respect to $u$.
Finally, we denote first and second derivative of the marginal density $f_{U}$ by $f_{U}^{\prime}$ and $f_{U}^{\prime\prime}$, respectively.
For a precise statement of our results we require the following assumptions.
Assumption 3.1.
$K(\cdot)$ is a symmetric density function supported on the interval (1,1)(-1,1), and $\int_{-1}^{1}K^{4}\left(u\right)\mathrm{d}u<\infty$; the kernel $L(\cdot)$ is the Langevin kernel defined in (2.10).
Assumption 3.2.
The density $f$ of $Y$ is twice differentiable.
  • (a)
    For any $u\in\mathbb{R}^{+}$, $\mathcal{D}_{f}^{2}(u,v)$ is uniformly continuous on $\mathbb{S}^{p-1}$.
  • (b)
    For any $v\in\mathbb{S}^{p-1}$, $\partial^{2}f(u,v)/\partial u^{2}$ is uniformly continuous on $\mathbb{R}^{+}$.
Proposition 3.1.
Suppose that Assumption 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$.
  • (i)
    If the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
    $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})=0.$ (3.5)
  • (ii)
    If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
    $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})$ $={\mathcal{M}}^{2}$ (3.6)
    $+\frac{h^{2}\phi_{2}(K)}{2}\left\{\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\left.\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{1})-\omega_{p-1}^{-1}\int_{\mathbb{R}^{+}}f_{U}(u)f_{U}^{\prime\prime}(u)\mathrm{d}u\right\}$ (3.7)
    $+\frac{\omega_{p-2}}{2\kappa}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})+O(\kappa^{-3/2}+h^{3}+h^{2}\kappa^{-1}),$ (3.8)
    where for $j\geq 1$
    $\phi_{j}(K)=\int_{-1}^{1}x^{j}K(u)du.$ (3.9)
By Proposition 3.1, the statistic $\hat{\mathcal{M}}^{2}_{n}$ is in general not an unbiased estimator of $\mathcal{M}^{2}$. The bias of the kernel estimation with respect to the radius $\|Y\|$ is of order $h^{2}$ and it is of order $1/\kappa$ with respect to the direction $Y/\|Y\|$. Therefore, smaller values of $h$ and larger values of $\kappa$ are preferred to diminish the bias the estimator $\hat{\mathcal{M}}_{n}^{2}$. We now turn to the weak convergence of $\hat{\mathcal{M}}_{n}^{2}$. Throughout this paper the symbol $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes weak convergence (convergence in distribution) of real valued random variables.
Theorem 3.1.
If 3.1 and 3.2 are satisfied, $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$, and the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
${\hat{\mathcal{M}}^{2}_{n}\over s_{n}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,1),$
where
$s_{n}^{2}=\frac{\psi_{2}(K)\kappa^{(p-1)/2}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)}{2^{p-2}\pi^{(p-1)/2}n(n-1)h},$ (3.10)
and
$\psi_{j}(K)=\int_{-1}^{1}K^{j}(u)du.$ (3.11)
Note that the convergence rate in Theorem 3.1 is of order $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, which is asymptotically greater than $\sqrt{n}$ as $\kappa^{(p-1)/2}/(nh)\to 0$, by assumption. On the other hand, weak convergence with a rate $\sqrt{n}$ is obtained in the case where the distribution of $Y$ is not spherical symmetric.
Theorem 3.2.
Assume that 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$. If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
$\sqrt{n}\big(\hat{\mathcal{M}}^{2}_{n}-\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]\big)\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2}),$
where the asymptotic variance is given by
$\sigma^{2}$ $=\int_{\mathbb{R}^{+}}\int_{\mathbb{S}^{p-1}}f(u,v)\big(f(u,v)-\omega_{p-1}^{-1}f_{U}(u)\big)^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)-({\mathcal{M}}^{2})^{2}.$ (3.12)
Remark 3.1.
Theorem 3.2 can be used to derive the weak convergence
$\sqrt{n}(\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2})$ (3.13)
in the case, where the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$. However, some care is necessary when replacing $\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]$ by $\mathcal{M}^{2}$ as the error of this replacement has to be of order $o(1/\sqrt{n})$. By Proposition 3.1(ii) this yields to the condition
$\sqrt{n}/\kappa+\sqrt{n}h^{2}\to 0,$ (3.14)
which has to be simultaneously satisfied with the condition $\kappa^{(p-1)/2}/(nh)\to 0$ required for controlling the variance of the $U$-statistic. This yields some restriction on the dimension $p$. For example, if $p=3$, the condition (3.14) can be satisfied by setting $h=n^{-3/8}$, $\kappa=\lfloor n^{9/16}\rfloor$. However, a simple calculation shows that (3.14) and $\kappa^{(p-1)/2}/(nh)\to 0$ cannot simultaneously hold if $p\geq 4$.
This issue can be mitigated via a simple bias reduction using the statistic
$\tilde{\mathcal{M}}^{2}_{n}(a,\kappa)=\frac{1}{1-a}\hat{\mathcal{M}}^{2}_{n}(\kappa)-\frac{a}{1-a}\hat{\mathcal{M}}^{2}_{n}(a\kappa),$ (3.15)
where $\hat{\mathcal{M}}^{2}_{n}(a\kappa)$ denotes the estimator $\hat{\mathcal{M}}^{2}_{n}$ using $a\kappa$ instead of $\kappa$ in the rapidly varying function $L(\cdot)$, see also Theorem 3 of Tsuruta, (2024). In addition, if we replace the kerne $K$ by its jackknife correction $\tilde{K}(x)=2\sqrt{2}K(\sqrt{2}x)-K(x)$, which satisfies $\phi_{2}(\tilde{K})=0$, we obtain from Proposition 3.1(ii)
$\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})$ $={\mathcal{M}}^{2}+O(\kappa^{-3/2}+h^{3}),$ (3.16)
and (3.13) holds under the conditions $\sqrt{n}/\kappa^{3/2}+\sqrt{n}h^{3}\to 0$ and $\kappa^{(p-1)/2}/(nh)\to 0$, which can be satisfied, whenever $p\leq 6$. If $p>6$ corresponding results hold under additional smoothness assumptions using the higher order expansions
$f(u,y)$ $=f(u,x)+\sum_{s=1}^{q}\frac{\theta^{s}}{s!}\xi^{\top}\mathcal{D}^{s}_{f}(u,x)\xi^{\otimes(s-1)}+O(\theta^{q+1}),$ (3.17)
$f(u_{0}+h,v)$ $=f(u_{0},v)+\sum_{s=1}^{q}\frac{h^{s}}{s!}\left.\frac{\partial^{s}f(u,v)}{\partial u^{s}}\right|_{u=u_{0}}+O(h^{q+1}),$ (3.18)
where $u^{\otimes s}$ denotes its $s$-th Kroneckerian power of the vector $u$. For example, if $q=3$, it follows observing $\int_{\Omega_{x}}\xi^{\top}\xi^{\otimes 2}\omega_{p-2}(\xi)=0$, and $\int x^{3}K(x)dx=0$ ($K$ is symmetry by Assumption 3.1) that the rate in (3.16) can be improved to $O(\kappa^{-2}+h^{4})$.
3.2 Some statistical applications
By Proposition 3.1, Theorem 3.1 and 3.2, the statistic $\hat{\mathcal{M}}_{n}^{2}$ is a consistent estimator of the measure $\mathcal{M}^{2}$, which defines the deviation from sphericity. To use these results for uncertainty quantification, one requires estimates of the variances $s_{n}^{2}$ and $\sigma^{2}$ defined in (3.10) and (3.12), respectively.
A simple estimator for $s_{n}^{2}$ is given by
$\hat{s}_{n}^{2}=\frac{\phi_{2}(K)\kappa^{(p-1)/2}}{2^{p-2}\pi^{(p-1)/2}n^{2}(n-1)^{2}c_{1}(\kappa)h^{2}}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j}),$ (3.19)
which is obtained by replacing integral of the squared density in (3.10) by its corresponding estimate (2.8). We will use this estimate to define a test for exact sphericity (see Remark 3.2 below).
The estimation of $\sigma^{2}$ is more difficult and we propose a jackknife approach for this purpose (see Chapter 5 of Lee,, 2003)). To be precise, let $\hat{\mathcal{M}}^{2}_{n-1}(-i)$ denote the estimator (2.12) of the minimum distance ${\cal M}^{2}$ based on the observations $Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}$, and define the pseudovalues
$\tilde{\mathcal{M}}^{2}_{i}=n\hat{\mathcal{M}_{n}}^{2}-(n-1)\hat{\mathcal{M}}^{2}_{n-1}(-i)$ (3.20)
($i=1,\ldots,n)$. The jackknife estimator of the asymptotic variance $\sigma^{2}=\lim_{n\to\infty}{\rm Var}(\sqrt{n}\mathcal{M}_{n}^{2})$ is then given by
$\hat{\sigma}^{2}_{n}=\frac{1}{4(n-1)}\sum_{i=1}^{n}(\tilde{\mathcal{M}}_{i}^{2}-\bar{\mathcal{M}}_{n}^{2})^{2},$ (3.21)
where
$\bar{\mathcal{M}}_{n}^{2}=n^{-1}\sum_{i=1}^{n}\tilde{\mathcal{M}}_{i}^{2}$ (3.22)
denotes the jackknife estimate of the mean of $\mathcal{M}_{n}^{2}$.
From Theorem 3.2 and Remark 3.1, we obtain a simple asymptotic confidence interval for $\mathcal{M}^{2}$, that is
$\hat{I}_{n}=\Big[\hat{\mathcal{M}}_{n}^{2}-{\hat{\sigma}_{n}\over\sqrt{n}}u_{1-\alpha/2},\hat{\mathcal{M}}_{n}^{2}+{\hat{\sigma}_{n}\over\sqrt{n}}u_{1-\alpha/2}\Big],$ (3.23)
where $u_{1-\alpha/2}$ denotes the $(1-\alpha/2)$-quantile of the standard normal distribution. The following result, which follows from Theorem 3.2 and the consistency of $\hat{\sigma}_{n}^{2}$ for $\sigma^{2}$, shows that this interval keeps its nominal level asymptotically.
Corollary 3.1.
If the assumptions in Theorem 3.2 and in Remark 3.1 are satisfied, we have
$\lim_{n\to\infty}\mathbb{P}\big(\mathcal{M}^{2}\in\hat{I}_{n}\big)=1-\alpha.$
Next, we turn to the problem of testing for spherical symmetry. As pointed out in the introduction we are not interested in testing for exact sphericity, that $\mathcal{M}^{2}=0$, because there are many applications where one does not really believe in exact sphericity, but wants to assume this with the hope that the deviations from sphericity are small and a procedure developed under the assumption of exact sphericity (such as classical ANOVA) still yields reliable and efficient inference. With this point of view, we propose to test the hypotheses
$H_{0}^{\rm rel}:\mathcal{M}^{2}\leq\Delta~~~~{\rm versus}~~~~H_{1}^{\rm rel}:\mathcal{M}^{2}>\Delta~,$ (3.24)
$H_{0}^{\rm eq}:\mathcal{M}^{2}\geq\Delta~~~~{\rm versus}~~~~H_{1}^{\rm eq}:\mathcal{M}^{2}<\Delta~,$ (3.25)
where $\Delta>0$ is a prespecified threshold. Note that this perspective of hypothesis testing aligns with the view expressed by Berger and Delampady, (1987), who argue that it is rare, and perhaps impossible, to have a null hypothesis that can be exactly modeled by a parameter being precisely 0. Similarly, Tukey, (1991), in the context of multiple comparisons of means, emphasizes that “All we know about the world teaches us that the effects of A and B are always different — in some decimal place — for any A and B. Thus, asking ‘Are the effects different?’ is foolish.” We call hypotheses the form (3.24) and (3.25) relevant hypotheses in the following discussion. Note also that hypotheses of the form (3.25) have found considerable attention in equivalence testing in the field of biostatistics, which explains the notations $H_{0}^{\rm eq}$ and $H_{1}^{\rm eq}$ in (3.25) (see Wellek,, 2010) and that rejection of the null hypothesis in (3.25) allows to decide for approximate sphericity at a controlled type I error.
We propose to reject the null hypothesis in (3.24) if
$\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}.$ (3.26)
Similarly, we propose to reject the null hypothesis in (3.25), whenever
$\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}.$ (3.27)
Our next results shows that both decision rules define consistent and asymptotic level $\alpha$ tests for the hypotheses (3.24) and (3.25), respectively.
Corollary 3.2.
Let the assumptions of Theorem 3.2 be satisfied.
  • (a)
    For the test (3.26), we have
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}>\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}<\Delta.\end{cases}$ (3.28)
  • (b)
    For the test (3.27), we have
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}<\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}>\Delta.\end{cases}$ (3.29)
We will illustrate the finite sample properties of the tests (3.26) and (3.27) in Section 5.
Remark 3.2.
  • (a)
    An important question in testing hypotheses of the form (3.24) and (3.25) is the choice of the threshold $\Delta$, a similar argument can be given for the hypotheses (3.24). Note that this can be done in a data adaptive way. To be precise we concentrate on the hypotheses (3.25), and note these hypotheses are nested, that is if $H_{0}^{\rm eq}$ is satisfied for $\Delta_{1}$, then it is also satisfied for all $\Delta_{2}\leq\Delta_{1}$. The decision rules for different $\Delta$’s are nested in the same (if $H_{0}^{\rm eq}$ is rejected for $\Delta_{2}$, it also rejected for all $\Delta_{1}\geq\Delta_{2}$). By the sequential rejection principle, we may simultaneously test the hypotheses in (3.25) for different $\Delta\geq 0$ starting at $\Delta=0$ and increasing $\Delta$ to find the minimum value of $\Delta$, say
    $\hat{\Delta}_{\alpha}:=\min\big\{\Delta\geq 0\,|\,\hat{\cal M}^{2}_{n}\leq\Delta+u_{\alpha}\hat{s}_{n}/\sqrt{n}\big\}$ (3.30)
    for which $H_{0}$ in (3.25) is rejected. The quantity $\hat{\Delta}_{\alpha}$ could be interpreted as a measure of evidence against the null hypothesis in (3.25). In this sense, the question of a reasonable choice of the threshold $\Delta$ can be postponed until after seeing the data.
  • (b)
    Although the focus of this paper is not on the classical hypothesis of exact sphericity, that this
    $H_{0}^{\rm exact}:\mathcal{M}^{2}=0~~~~\text{versus}~~~~\mathcal{H}_{1}^{\rm exact}:\mathcal{M}^{2}>0~,$ (3.31)
    it is worthwhile to mention that the theory developed so far also provides a test for these hypotheses. More precisely, we propose to reject the null hypothesis in (3.31), whenever
    $\hat{\mathcal{M}}_{n}^{2}>u_{1-\alpha}{\hat{s}_{n}\over\sqrt{n}},$ (3.32)
    where $\hat{s}_{n}$ is defined in (3.19). It then follows from Theorem 3.1 and 3.2 that this decision rule defines a consistent and asymptotic level $\alpha$ test.
4 Pivotal inference
The statistical methodology developed in Section 3.2 requires estimation of the variance $\hat{s}_{n}^{2}$. In this section, we develop a pivotal confidence interval for the deviation $\mathcal{M}^{2}$ from sphericity and pivotal tests for the hypotheses (3.24) and (3.25). For this purpose, we prove a weak convergence result for a sequential version of the estimator $\hat{\mathcal{M}}_{n}^{2}$.
To be precise, let $\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}$ denote the estimator (2.12) calculated for the sample $Y_{1},\ldots,Y_{\lfloor nt\rfloor}$, where $t\in[0,1]$. We consider the sequential process
$S_{n}(t)=\frac{\lfloor nt\rfloor}{2\sqrt{n}}\big(\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}-\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]\big),$ (4.1)
whose asymptotic properties are investigated in following theorem. Throughout this section the symbol $\Rightarrow$ denotes weak convergence in the space $\ell^{\infty}([0,1])$ of bounded functions on the interval $[0,1]$.
Theorem 4.1.
If the conditions of Theorem 3.2 are satisfied and $(\log n)^{2}\kappa^{(p-1)/2}/(nh)\to 0$, we have
$\big\{S_{n}(t)\big\}_{t\in[0,1]}\Rightarrow\big\{\sigma\mathbb{B}(t)\big\}_{t\in[0,1]},$ (4.2)
where $\big\{\mathbb{B}(t)\big\}_{t\in[0,1]}$ is a standard Brownian motion and $\sigma^{2}$ is defined in (3.12).
Remark 4.1.
It follows from Proposition 3.1 that
$\frac{1}{\sqrt{n}}\max_{2\leq k\leq n}\Big|k\big\{\mathcal{M}^{2}-\mathbb{E}[H_{n}(Y_{1},Y_{2})]\big\}\Big|=\sqrt{n}\big|\mathcal{M}^{2}-\mathbb{E}[H_{n}(Y_{1},Y_{2})]\big|=O(\sqrt{n}/\kappa+\sqrt{n}h^{2})=o(1)$
provided that the bandwidth condition (3.14) is satisfied. Along with Theorem 4.1, we therefore have
$\frac{\lfloor nt\rfloor}{2\sqrt{n}}\big(\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}-{\mathcal{M}}^{2})\Rightarrow\big\{\sigma\mathbb{B}(t)\big\}_{t\in[0,1]}.$ (4.3)
The bandwidth condition (3.14) can be relaxed as discussed in Remark 3.1.
We will now use the result (4.3) to develop pivotal inference for the measure $\mathcal{M}^{2}$. To be precise, we define the statistic
$\hat{V}_{n}=\int^{1}_{0}\big|\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}-\hat{\mathcal{M}}^{2}_{n}\big|\,t\,\mathrm{d}t,$ (4.4)
and note that it follows from the continuous mapping theorem and Theorem 4.1 that
$\frac{\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2}}{\hat{V}_{n}}\;\stackrel{{\scriptstyle d}}{{\longrightarrow}}\;W=\frac{\mathbb{B}(1)}{\int^{1}_{0}|\,\mathbb{B}(t)-t\mathbb{B}(1)|\mathrm{d}t}.$ (4.5)
In the following, let $q_{1-\alpha}$ denote the $(1-\alpha)$ quantile of the distribution of $W$. Then, a pivotal confidence interval for $\mathcal{M}^{2}$ is given by
$\hat{I}_{n}^{\rm piv}=\Big[\hat{\mathcal{M}}_{n}^{2}-\hat{V}_{n}q_{1-\alpha/2},\hat{\mathcal{M}}_{n}^{2}+\hat{V}_{n}q_{1-\alpha/2}\Big],$ (4.6)
Similarly, we propose as a test for the hypotheses (3.24) and (3.25) to reject the null hypothesis, whenever
$\hat{\mathcal{M}}_{n}^{2}$ $>\Delta+q_{1-\alpha}{\hat{V}_{n}},$ (4.7)
and
$\hat{\mathcal{M}}_{n}^{2}$ $\leq\Delta+q_{\alpha}{\hat{V}_{n}},$ (4.8)
respectively. We note that the comments regarding the choice of the threshold $\Delta$ in Remark 3.2(a) remain valid for the pivotal tests (4.7) and (4.8), and summarize the properties of these tests in the following corollary.
Corollary 4.1.
If the assumptions in Theorem 4.1 and in Remark 4.1 are satisfied, we have
$\lim_{n\to\infty}\mathbb{P}\big(\mathcal{M}^{2}\in\hat{I}_{n}^{\rm piv}\big)=1-\alpha.$
Moreover, the tests (4.7) and (4.8) are consistent and have asymptotic level $\alpha$ for the hypotheses (3.24) and (3.25), respectively.
5 Finite sample properties
In this section we illustrate the finite sample properties of the developed methodology by means of a small simulation study. All results presented here are based on $1000$ simulation runs.
According to Theorem 3.2, the term $\kappa^{(p-1)/2}/(nh)$ controls the variance for the non-leading term in the Gaussian approximation, while $\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}$ accounts for the bias term with the bias reduction procedure discussed in Remark 3.2. In addition, for the sequential convergence, by Theorem 4.1, the inflated rate $(\log n)^{2}\kappa^{(p-1)/2}/(nh)$ controls the variance of the process of the non-leading term. In order to approximate Gaussianity in finite samples better, we recommend imposing a stronger control over the non-leading term, that is, $n^{3/2}\kappa^{(p-1)/2}/(nh)\to 0$. Therefore, we consider bandwidths $(h,\kappa)$ satisfying
$(\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4})^{2}\asymp\kappa^{(p-1)/2}\sqrt{n}/h.$ (5.1)
For a data-dependent choice of the smoothing parameters, we first define a sequence for $(h,\kappa)$’s, i.e., $(h_{i},\kappa_{i}):=(n^{-1/(2(p+8))}a_{i},n^{1/(p+8)}c_{i})$ satisfying (5.1), where $a_{i}$ and $c_{i}$ are prespecified constants, $i=1,\ldots,M$. The choices of $a_{i}$ and $c_{i}$ can be obtained by inspecting the turning point of the graph of $\hat{\mathcal{M}}_{n}^{2}$ versus $h$ and that versus $\kappa$. We give more details for each model below.
Given the chosen grid, we select the pair of parameters that minimizes the volatility of the self-normalizing term $\hat{V}_{n}$ in (4.4), since its expectation is proportional to the standard deviation of the statistic $\hat{\cal M}_{n}^{2}$. More precisely, we select $(h,\kappa)$ as the minimizer of
$\min_{i\in\{1,\ldots,M\}}\mathrm{SE}\{\hat{V}_{n}(h_{j},\kappa_{j})\}_{j=i-1}^{i+1},$ (5.2)
where SE denotes the standard error.
In the following discussion we consider two models.
Model 1: $Y$ has a $3$-dimensional Gaussian distribution with mean vector $\mu=(1,0,2)^{\top}$ and covariance matrix
$\Sigma=0.25\times\begin{pmatrix}1&0.3&0\\ 0.3&1&0\\ 0&0&1\end{pmatrix}$. (5.3)
The deviation $\mathcal{M}^{2}$ was calculated numerically as $\mathcal{M}^{2}\approx 0.95$ , using a using a larger sample size $2500$ and averaged over $200000$ times of repeated simulation. For the selection of the smoothing parameters via (5.2), we use
$(c_{i})_{i=1}^{5}=(72.50,73.75,75.00,76.25,77.50)^{\top}~,~~(a_{i})_{i=1}^{5}=(0.7500,0.8125,0.8750,0.9375,1.0000)^{\top}.$
Model 2: $Y$ has a $5$-dimensional Gaussian distribution with mean vector $\mu=(1,0,0,-2,0)^{\top}$ and covariance matrix
$\Sigma=0.25\times\begin{pmatrix}1&0.2&0&0&0\\ 0.2&1&0.3&0&0\\ 0&0.3&1&0&0\\ 0&0&0&1&0.2\\ 0&0&0&0.2&1\end{pmatrix}$. (5.4)
The deviation $\mathcal{M}^{2}$ was calculated numerically as $\mathcal{M}^{2}\approx 1.97$. For the selection of the smoothing parameters via (5.2), we use
$(c_{i})_{i=1}^{5}=(37.50,38.75,40.00,41.25,42.50)^{\top}~,~~(a_{i})_{i=1}^{5}=(0.7500,0.8125,0.8750,0.9375,1.0000)^{\top}.$
5.1 Confidence intervals for the deviation from a spherical distribution
In Table 1 and 2, we display the empirical coverage rates and the average widths of Jackknife confidence intervals (3.23) and pivotal confidence intervals (4.6) for the minimum distance (2.4), where we choose the nominal levels as $95\%$ and $90\%$.
Jackknife Pivotal
coverage width coverage width
$n$ 95% 90% 95% 90% 95% 90% 95% 90%
200 95.9 91.7 0.46 0.39 97.8 95.3 0.71 0.56
300 96.5 91.7 0.36 0.30 97.4 93.7 0.53 0.42
400 94.6 89.6 0.31 0.26 96.7 92.0 0.45 0.35
500 95.4 91.0 0.27 0.23 97.0 92.5 0.39 0.31
600 94.9 89.8 0.25 0.21 96.5 91.2 0.36 0.28
800 94.5 90.8 0.21 0.18 95.2 90.4 0.30 0.23
1000 95.4 89.1 0.19 0.16 95.2 89.8 0.25 0.20
Table 1: Simulated coverage rates (in %) and average widths of Jackknife and pivotal confidence intervals for Model 1.
Jackknife Pivotal
coverage width coverage width
$n$ 95% 90% 95% 90% 95% 90% 95% 90%
200 94.7 90.1 1.19 1.00 98.1 94.9 1.87 1.46
300 95.0 91.0 0.94 0.79 97.0 93.0 1.42 1.11
400 94.8 90.3 0.80 0.67 97.6 93.8 1.17 0.92
500 94.1 89.4 0.71 0.59 96.3 92.8 1.02 0.80
600 93.9 90.0 0.65 0.54 95.9 91.4 0.93 0.72
800 93.5 88.1 0.55 0.46 95.0 90.3 0.78 0.61
1000 94.3 89.4 0.49 0.41 94.7 90.4 0.69 0.54
Table 2: Simulated coverage rates (in %) and average widths of the Jackknife and pivotal confidence intervals for Model 2.
For Model 1, the empirical coverage rates of both methods are close to their nominal levels for sample sizes larger than $500$. The Jackknife confidence interval achieves more accurate simulated coverage rates when the sample size is small, while the pivotal confidence intervals are more conservative for smaller sample sizes. A potential explanation of this observation is that in this case the widths of the pivotal method are much larger than those of the Jackknife method.
We observe similar patterns for Model 2. The pivotal confidence intervals are more conservative compared to the Jackknife confidence intervals but achieve accurate coverage rates for large sample sizes. At the same time, the jackknife confidence intervals have narrower widths and work very well even when the sample size is small.
5.2 Relevant hypothesis
In Table 3 and 4, we display the simulated rejection rates of Jackknife test (3.27) and the pivotal test for the hypotheses (3.25) for different values of $\Delta$ and a significance level of $5\%$. For both models the results reflect the asymptotic properties of the test described in Corollary 3.2. At the “boundary” of the hypotheses, where ${\cal M}^{2}=\Delta$, the empirical sizes of the pivotal test (4.8) are close to the nominal level $5\%$ in most cases, while the type I error of the Jackknife test (3.27) is too large for small sample sizes.
$H_{0}$ $H_{1}$
$\Delta=0.7$ $\Delta=0.9$ $\Delta=\mathcal{M}^{2}$ $\Delta=1.1$ $\Delta=1.3$ $\Delta=1.5$
$n$ (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8)
200 0 0 2.0 1.4 7.7 3.9 39.2 23.1 87.8 68.2 99.1 91.3
300 0 0 1.3 0.8 6.1 4.2 51.9 36.2 95.4 82.4 100.0 98.3
400 0 0 0.9 0.8 5.8 4.9 60.3 43.0 98.7 91.5 100.0 99.7
500 0 0 0.5 0.6 6.2 4.5 71.2 53.7 99.6 95.9 100.0 100.0
600 0 0 0.5 0.7 5.1 3.9 74.3 59.6 99.8 97.6 100.0 99.9
800 0 0 0.5 0.8 3.9 3.9 83.0 65.6 100.0 99.3 100.0 100.0
1000 0 0 0.2 0.2 5.6 4.9 89.8 77.0 100.0 99.6 100.0 100.0
Table 3: Simulated rejection rates (in %) of the Jackknife test (3.27) and pivotal test (4.8) for the relevant hypotheses (3.25) in Model 1.
In the “interior of the null hypothesis”, where $\mathcal{M}^{2}>\Delta$, the rejection rates of both tests decreases to 0 as $\Delta$ decreases and the sample size increases. Under the alternative, where $\mathcal{M}^{2}<\Delta$, the rejection rates of both tests increase to $1$ as $\Delta$ increases and the sample size increases. The rejection rates of the Jackknife test (3.27) are usually higher than those of the pivotal test (4.8).
$H_{0}$ $H_{1}$
$\Delta=1.7$ $\Delta=1.8$ $\Delta=\mathcal{M}^{2}$ $\Delta=2.3$ $\Delta=2.5$ $\Delta=2.8$
$n$ (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8)
200 0.8 0.2 3.2 1.3 8.0 4.5 36.1 24.2 60.8 43.0 85.3 69.1
300 0.2 0.3 1.7 1.1 8.5 4.4 46.8 32.1 72.9 55.5 93.6 81.0
400 0.0 0.0 1.7 1.4 6.0 3.8 57.3 41.7 82.8 67.5 97.4 90.9
500 0.2 0.1 1.0 0.7 6.9 4.6 63.3 46.3 90.6 76.7 99.5 94.4
600 0.1 0.1 0.8 1.1 5.9 4.8 66.7 51.9 93.2 81.4 99.7 96.4
800 0.0 0.0 0.1 0.3 5.3 3.8 78.4 59.9 97.7 89.0 100.0 98.7
1000 0.0 0.0 0.2 0.2 4.9 4.8 82.8 71.0 98.7 92.6 100.0 99.3
Table 4: Simulated rejection rates (in %) of Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in Model 2.
Summarizing, for testing for relevant deviations from sphericity from independent identically distributed data, the Jackknife test exhibits some advantages compared to the pivotal test if the sample size is sufficiently large. However, we emphasize that this observation can only be made for independent data. In the case of dependencies, the Jackknife estimator does not yield a valid testing procedure. On the other hand, a careful inspection of the proofs in the appendix shows that the asymptotic statements in Section 4 remain valid for stationary processes under appropriate mixing (Bradley,, 2007), physical dependence (Wu,, 2005) or $m$-approximability (Hörmann and Kokoszka,, 2010) conditions. Consequently, Corollary 4.1 remains valid as well, and the pivotal test (4.8) has asymptotic level $\alpha$ and is consistent for the hypotheses (3.25).
We illustrate this fact by a small simulation for the tests (3.27) and (4.8) in a model with dependent data. To be precise, let $Z_{i}=(z_{i,1},z_{i,2},z_{i,3})^{\top}$, where
$z_{i,j}=0.3z_{i-i,j}+\epsilon_{i,j},\quad j=1,2,3,$ (5.5)
and the $\epsilon_{i,j}^{\prime}s$ are independent standard normal distributed random variables. The data $(Y_{i})_{i=1}^{n}$ is then generated by
$Y_{i}=\Sigma^{1/2}Z_{i}\times\sqrt{1-0.3^{2}}+\mu,$ (5.6)
where $\Sigma$ and $\mu$ are defined in Model 1. We display in Table 5 the rejection probabilities of both tests for dependent data. In the interior of the null hypothesis and under the alternative we observe a quantitative similar behavior as for independent data. However, at the boundary ($\mathcal{M}^{2}=\Delta$) the Jackknife test (3.27) does not keep its nominal level $5\%$ (in all considered cases). On the other hand, the pivotal test (4.8) yields a very good approximation of the nominal level.
$H_{0}$ $H_{1}$
$\Delta=0.7$ $\Delta=0.9$ $\Delta=\mathcal{M}^{2}$ $\Delta=1.1$ $\Delta=1.3$ $\Delta=1.5$
$n$ (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8) (3.27) (4.8)
200 0 0 2.2 0.8 7.0 3.5 38.1 23.4 86.4 63.1 98.6 89.7
300 0 0 2.4 1.1 9.0 4.7 48.2 30.7 93.3 81.2 99.9 97.1
400 0 0 1.3 0.9 8.2 4.3 55.9 39.8 97.3 87.5 100.0 98.6
500 0 0 1.1 1.2 7.6 4.8 64.7 44.8 99.6 92.8 100.0 99.6
600 0 0 1.4 0.4 10.1 5.4 70.1 52.1 99.8 94.7 100.0 99.7
800 0 0 0.9 0.5 7.8 5.0 79.9 59.2 100.0 98.1 100.0 100.0
1000 0 0 0.9 0.4 8.4 5.4 86.0 66.8 100.0 99.5 100.0 100.0
Table 5: Simulated rejection rates (in %) of the Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in the case of dependent data.
5.3 Data analysis
For illustrate the potential applications of our approach , we analyze the log-returns of the daily exchange rate (close price) of the Yen to the Dollar and the Pound to the Euro from January 2, 2009, to December 31, 2009, which has also been investigated in Einmahl and Gantner, (2012) using a test for independent data. We downloaded the data from Yahoo Finance via the R package “quantmod” and obtained 260 entries for the log-returns, as the transactions of the currency can be conducted on weekdays including holidays. The mean of the bivariate series turns out to be $9.4\times 10^{-5}$ and $2.7\times 10^{-4}$. In Figure 1, we plot the absolute autocorrelations of different lags of both of the series. We found that there exists some evidence of serial dependence, for example, the lag 12 of the Yen to Dollar and the lag 6 of Pound to Euro.
Figure 1: Absolute autocorrelations of the log-returns of the daily exchange rate of the Yen to the Dollar and the Pound to the Euro from January 2nd, 2009, to December 31st, 2009.
[…]
Acknowledgements This work was partially supported by DFG Research unit 5381 Mathematical Statistics in the Information Age (Project number 460867398) and by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
Appendix A Preliminary technical results
In this section we present and prove several preliminary results, which will be essential for the proofs of the statements in Section 3 and 4. Our first result provides some properties of the Fisher-von-Mises distribution.
Lemma A.1.
For the Fisher-von-Mises distribution in (2.10) we have for any $j\geq 1$, $p\geq 2$
$b_{j}(\kappa)$ $=\omega_{p-2}\int_{0}^{\pi}L(\kappa\cos\theta)(\sin\theta)^{p-2}\theta^{j}\mathrm{d}\theta\sim a_{j}(p)\kappa^{-j/2}$ (5.7)
as $\kappa\to\infty$, where $a_{j}(p)=2^{j/2}\Gamma((p+j-1)/2)/\Gamma((p-1)/2)$. Similarly,
$c_{j}(\kappa)$ $=\omega_{p-2}\int_{0}^{\pi}L^{j}(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta\sim d_{j}(p)\kappa^{(j-1)(p-1)/2}$ (5.8)
as $\kappa\to\infty$, where $d_{j}(p)=2^{(1-j)(p-1)/2}j^{-(p-1)/2}\pi^{(1-j)(p-1)/2}$. Moreover, for any $t>0$ we have
${L^{j}(\kappa t)\over L(j\kappa t)}\sim\kappa^{(j-1)(p-1)/2}j^{-p/2+1/2}(2\pi)^{(j-1)(p-1)/2}$ (5.9)
as $\kappa\to\infty$.
Proof.
By equation (9) in Marco Di Marzio and Taylor, (2014), we have
$b_{j}(\kappa)\sim\frac{2^{j/2}\Gamma((p+j-1)/2)}{\kappa^{j/2}\Gamma((p-1)/2)}$ (5.10)
as $\kappa\to\infty$ and with the notation $a_{j}(p)=2^{j/2}\Gamma((p+j-1)/2)/\Gamma((p-1)/2)$ the result (5.7) follows. For a proof of (5.8) we note that the modified Bessel function in (2.11) satisfies $\mathcal{I}_{v}(\kappa)\sim e^{\kappa}/\sqrt{2\pi\kappa}$, as $\kappa\to\infty$ (see equation (9.6.18) on Page 376 of Abramowitz and Stegun, (1968)). Therefore, recalling the definition of the Fisher-von-Mises distribution in (2.10), we obtain
$L(\kappa t)\sim\kappa^{p/2-1}\{(2\pi)^{p/2}e^{\kappa}/\sqrt{2\pi\kappa}\}^{-1}e^{\kappa t},$ (5.11)
which yields (with the substitution $t=\cos\theta$)
$c_{j}(\kappa)$ $=\omega_{p-2}\int_{-1}^{1}L^{j}(\kappa t)(1-t^{2})^{(p-3)/2}\mathrm{d}t$ (5.12)
$\sim\omega_{p-2}\kappa^{jp/2-j/2}\{(2\pi)^{(p-1)/2}e^{k}\}^{-j}\int_{-1}^{1}e^{j\kappa t}(1-t^{2})^{(p-3)/2}\mathrm{d}t.$ (5.13)
Observing that
$\int_{-1}^{1}e^{j\kappa t}(1-t^{2})^{(p-3)/2}\mathrm{d}t$~ $=\int_{0}^{2}e^{j\kappa(1-t)}\{t(2-t)\}^{(p-3)/2}\mathrm{d}t$ (5.14)
$=\int_{0}^{2}e^{j\kappa}e^{-j\kappa t}t^{(p-3)/2}(2-t)^{(p-3)/2}\mathrm{d}t$ (5.15)
$\sim 2^{(p-3)/2}e^{j\kappa}(j\kappa)^{-(p-1)/2}\int_{0}^{2j\kappa}e^{-t}t^{(p-3)/2}\mathrm{d}t$ (5.16)
$\sim 2^{(p-3)/2}e^{j\kappa}(j\kappa)^{-(p-1)/2}\Gamma\{(p-1)/2\},$ (5.17)
we have for $j\geq 1$
$c_{j}(\kappa)$ $\sim\omega_{p-2}(\kappa/(2\pi))^{pj/2-j/2}2^{(p-3)/2}(j\kappa)^{-(p-1)/2}\Gamma\{(p-1)/2\}$ (5.18)
$\sim\kappa^{(j-1)(p-1)/2}2^{(1-j)(p-1)/2}j^{-(p-1)/2}\pi^{(1-j)(p-1)/2},$ (5.19)
where we used $\omega_{p-2}=2\pi^{(p-1)/2}/\Gamma((p-1)/2)$ in the last step. The assertion (5.8) now follows with the notation $d_{j}(p)=2^{(1-j)(p-1)/2}j^{-(p-1)/2}\pi^{(1-j)(p-1)/2}$. Finally, (5.9) is obtained by a direct calculation using the expansion (5.11). ∎
We will use Lemma A.1 for the calculation of the moments of the statistic (2.12), which is complicated due to the definition of the kernel $H_{n}$. We present two auxiliary results, which are required as intermediate steps in these calculations.
Lemma A.2.
If Assumption 3.1 and 3.2 are satisfied, we have
$h^{-1}c_{1}^{-1}(\kappa)\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ $=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.20)
$+\frac{h^{2}\phi_{2}(K)}{2}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left.f(u_{2},v_{1})\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.21)
$+\frac{\omega_{p-2}b_{2}(\kappa)}{2c_{1}(\kappa)(p-1)}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.22)
$+O(c_{1}^{-1}(\kappa)b_{3}(\kappa)+b_{2}(\kappa)h^{2}+h^{3}),$ (5.23)
$h^{-1}c_{1}^{-2}(\kappa)\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{2}(\kappa V_{i}^{\top}V_{j})\Big]$ $=c_{1}^{-2}(\kappa)c_{2}(\kappa)\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\mathrm{d}v$ (5.24)
$+O(c_{1}^{-2}(\kappa)c_{2}(\kappa)h^{2}+c_{1}^{-2}(\kappa)\kappa^{(p-1)/2}b_{2}(2\kappa)),$ (5.25)
$h^{-1}c_{1}^{-1}(\kappa)\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ $=\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\mathrm{d}v+O(h^{2}+c_{1}^{-1}(\kappa)b_{2}(\kappa)),$ (5.26)
$h^{-1}\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $=\int_{\mathbb{R}^{+}}f_{U}^{2}(u)du+\frac{h^{2}\phi_{2}(K)}{2}\int_{\mathbb{R}^{+}}f_{U}(u)f^{\prime\prime}_{U}(u)du+O(h^{3}),$ (5.27)
$h^{-1}\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $=\psi_{2}(K)\int_{\mathbb{R}^{+}}f_{U}^{2}(u)du+O(h^{2}),$ (5.28)
where $\psi_{j}(K)=\int_{-1}^{1}K^{j}(u)\mathrm{d}u$, $\phi_{j}(K)=\int_{-1}^{1}x^{j}K(u)\mathrm{d}u$.
Proof.
$\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ (5.29)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\int_{\mathbb{R}^{+}}K\left(\frac{u_{1}-u_{2}}{h}\right)L(\kappa v_{1}^{\top}v_{2})f(u_{1},v_{1})f(u_{2},v_{2})\mathrm{d}u_{1}\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{1})\ \omega_{p-1}(\mathrm{d}v_{2}).$ (5.30)
By a change of variables, that is $u=(u_{1}-u_{2})/h$, $\cos\theta=v_{1}^{\top}v_{2}$, we obtain
$u_{1}=hu+u_{2},\quad v_{2}=v_{1}\cos\theta+\xi\sin\theta,\quad\xi\perp v_{1}.$ (5.31)
Note that $\xi\in\Omega_{v_{1}}:=\{\xi\in\mathbb{S}^{p-1}:~\xi\perp v_{2}\}$, so that
$\omega_{p-1}(dv_{2})=(\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi).$ (5.32)
Then, with (5.31) and (5.32), we can rewrite (5.30) into
$\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]=h\int_{\Omega_{v_{1}}}\int_{0}^{\pi}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\int_{\mathbb{R}^{+}}K\left(u\right)L(\kappa\cos\theta)f(hu+u_{2},v_{1})$ (5.33)
$\times f(u_{2},v_{1}\cos\theta+\xi\sin\theta)\mathrm{d}u\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})\ (\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi).$ (5.34)
Recall the notations
$b_{j}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L(\kappa\cos\theta)(\sin\theta)^{p-2}\theta^{j}\mathrm{d}\theta,$ (5.35)
$c_{j}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L^{j}(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta$ (5.36)
and note that (see the proof of Theorem 3.1 in Marco Di Marzio and Taylor, (2014))
$\int_{\Omega_{v_{1}}}\xi\omega_{p-2}(\mathrm{d}\xi)=\mathbf{0}_{p},\quad\int_{\Omega_{v_{1}}}\xi\xi^{\top}\omega_{p-2}(\mathrm{d}\xi)={\omega_{p-2}\over(p-1)}(\mathbf{I}_{p}-v_{1}v_{1}^{\top}),$ (5.37)
where $\mathbf{0}_{p}$ and $\mathbf{I}_{p}$ denote the $p$-dimensional origin and $p\times p$ identity matrix, respectively. Moreover, $f(u,v_{1})$ is defined on $\mathbb{R}^{+}\times\mathbb{S}^{p-1}$ and
$v_{1}^{\top}D^{2}_{f}(u,v_{1})v_{1}=0$ (5.38)
(see the proof of Theorem 3.1 in Marco Di Marzio and Taylor,, 2014), which yields
$\int_{\Omega_{v_{1}}}\xi^{\top}D_{f}(u_{2},v_{1})\omega_{p-2}(\mathrm{d}\xi)$ $=0,$ (5.39)
$\int_{\Omega_{v_{1}}}\xi^{\top}D^{2}_{f}(u_{2},v_{1})\xi\ \omega_{p-2}(\mathrm{d}\xi)$ $=\int_{\Omega_{v_{1}}}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\xi\xi^{\top}\}\ \omega_{p-2}(\mathrm{d}\xi)$ (5.40)
$=\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})(\omega_{p-2}(p-1)^{-1}(\mathbf{I}_{p}-v_{1}v_{1}^{\top}))\}$ (5.41)
$=\omega_{p-2}(p-1)^{-1}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}.$ (5.42)
Using the Taylor expansions in (3.4) and (3.1), we therefore obtain for (5.34)
$h^{-1}\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ (5.43)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\int_{-1}^{1}K\left(u\right)[f(u_{2},v_{1})+hu\left.\frac{\partial f(u,v_{1})}{\partial u}\right|_{u=u_{2}}+\frac{h^{2}u^{2}}{2}\left.\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}+O(h^{3})]du\right\}$ (5.44)
$\times\left\{\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)[f(u_{2},v_{1})+\theta\xi^{\top}D_{f}(u_{2},v_{1})\right$. (5.45)
$~~~~~~~~\left.+\frac{\theta^{2}}{2}\xi^{\top}D^{2}_{f}(u_{2},v_{1})\xi+O(\theta^{3})]\ (\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.46)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{f(u_{2},v_{1})+\frac{h^{2}\phi_{2}(K)}{2}\left.\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}+O(h^{3})\right\}$ (5.47)
$\times\left\{c_{1}(\kappa)f(u_{2},v_{1})+\frac{\omega_{p-2}}{2(p-1)}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)\theta^{2}(\sin\theta)^{p-2}\mathrm{d}\theta\right$. (5.48)
$~~~~~~~~\left.+\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)O(\theta^{3})(\sin\theta)^{p-2}\mathrm{d}\theta\omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.49)
$=c_{1}(\kappa)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u_{2},v_{1})\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.50)
$~~~~~~~~+c_{1}(\kappa)\frac{h^{2}\phi_{2}(K)}{2}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left.f(u_{2},v_{1})\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.51)
$~~~~~~~~+\frac{\omega_{p-2}b_{2}(\kappa)}{2(p-1)}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})+O(b_{3}(\kappa)+h^{3}+b_{2}(\kappa)h^{2}),$ (5.52)
where we have used $\int_{-1}^{1}uK(u)\mathrm{d}u=0$, (5.39) and (5.42) in the second equality.
Recalling the notation
$c_{2}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L^{2}(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta,$ (5.53)
it follows by similar arguments and Lemma A.1 that
$h^{-1}\mathbb{E}\Big[K^{2}\left(\frac{U_{i}-U_{j}}{h}\right)L^{2}(\kappa V_{i}^{\top}V_{j})\Big]$ (5.54)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\int_{-1}^{1}K^{2}\left(u\right)[f(u_{2},v_{1})+hu\left.\frac{\partial f(u,v)}{\partial u}\right|_{u=u_{2}}+O(h^{2})]du\right\}$ (5.55)
$\times\left\{\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L^{2}\left(\kappa\cos\theta\right)[f(u_{2},v_{1})+\theta\xi^{\top}D_{f}(u_{2},v_{1})\right$. (5.56)
$~~~~~~~\left.+\frac{\theta^{2}}{2}\xi^{\top}D^{2}_{f}(u_{2},v_{1})\xi+O(\theta^{3})]\ (\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.57)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\psi_{2}(K)f(u_{2},v_{1})+O(h^{2})\right\}\times\left\{c_{2}(\kappa)f(u_{2},v_{1})\right$. (5.58)
$~~~~~~~\left.+\frac{\omega_{p-2}}{2(p-1)}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\int_{0}^{\pi}L^{2}\left(\kappa\cos\theta\right)\theta^{2}(\sin\theta)^{p-2}\mathrm{d}\theta\right$. (5.59)
$~~~~~~~\left.+\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L^{2}\left(\kappa\cos\theta\right)O(\theta^{3})(\sin\theta)^{p-2}\mathrm{d}\theta\omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.60)
$=c_{2}(\kappa)\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u_{2},v_{1})\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})+O(c_{2}(\kappa)h^{2}+\kappa^{(p-1)/2}b_{2}(2\kappa)).$ (5.61)
and
$h^{-1}\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ (5.62)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\int_{-1}^{1}K^{2}\left(u\right)[f(u_{2},v_{1})+hu\left.\frac{\partial f(u,v)}{\partial u}\right|_{u=u_{2}}+O(h^{2})]du\right\}$ (5.63)
$\times\left\{\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)[f(u_{2},v_{1})+\theta\xi^{\top}D_{f}(u_{2},v_{1})\right$. (5.64)
$~~~~~~~\left.+\frac{\theta^{2}}{2}\xi^{\top}D^{2}_{f}(u_{2},v_{1})\xi+O(\theta^{3})]\ (\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.65)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\psi_{2}(K)f(u_{2},v_{1})+O(h^{2})\right\}\times\left\{c_{1}(\kappa)f(u_{2},v_{1})\right$. (5.66)
$~~~~~~~\left.+\frac{\omega_{p-2}}{2(p-1)}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)\theta^{2}(\sin\theta)^{p-2}\mathrm{d}\theta\right$. (5.67)
$~~~~~~~\left.+\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L\left(\kappa\cos\theta\right)O(\theta^{3})(\sin\theta)^{p-2}\mathrm{d}\theta\omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})$ (5.68)
$=c_{1}(\kappa)\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u_{2},v_{1})\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})+O(c_{1}(\kappa)h^{2}+b_{2}(\kappa)).$ (5.69)
Finally, with simpler arguments, we can also derive
$h^{-1}\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $=\int_{\mathbb{R}^{+}}f_{U}^{2}(u)du+\frac{h^{2}\phi_{2}(K)}{2}\int_{\mathbb{R}^{+}}f_{U}(u)f^{\prime\prime}_{U}(u)du+O(h^{3}).$ (5.70)
$h^{-1}\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $=\psi_{2}(K)\int_{\mathbb{R}^{+}}f_{U}^{2}(u)du+O(h^{2}),$ (5.71)
which completes the proof of Lemma A.2. ∎
Lemma A.3.
If Assumption 3.1 and 3.2 are satisfied, we have
$h^{-1}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{j}(\kappa V_{i}^{\top}V_{j})\ \Big]$ $\sim\psi_{4}(K)c_{j}(\kappa)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.72)
$h^{-1}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $\sim\psi_{4}(K)\int_{\mathbb{R}^{+}}f_{U}^{2}(u)\mathrm{d}u.$ (5.73)
Proof.
Recall the notations
$\psi_{j}(K)=\int_{-1}^{1}K^{j}\left(u\right)\mathrm{d}u,\quad c_{j}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L^{j}(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta.$ (5.74)
and the Taylor expansions in (3.4) and (3.1), by (5.39) and (5.42), it follows
$h^{-1}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{j}(\kappa V_{i}^{\top}V_{j})\Big]$ (5.75)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\int_{-1}^{1}K^{4}\left(u\right)[f(u_{2},v_{1})+hu\left.\frac{\partial f(u,v)}{\partial u}\right|_{u=u_{2}}+O(h^{2})]du\right\}$ (5.76)
$\times\left\{\int_{\Omega_{v_{1}}}\int_{0}^{\pi}L^{j}\left(\kappa\cos\theta\right)[f(u_{2},v_{1})+\theta\xi^{\top}D_{f}(u_{2},v_{1})+\frac{\theta^{2}}{2}\xi^{\top}D^{2}_{f}(u_{2},v_{1})\xi+O(\theta^{3})]\right$. (5.77)
$\left.\times(\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{1})$ (5.78)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{\psi_{4}(K)f(u_{2},v_{1})+O(h^{2})\right\}\times\left\{c_{j}(\kappa)f(u_{2},v_{1})\right$. (5.79)
$~~~~~~~\left.+\frac{\omega_{p-2}}{2(p-1)}\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\int_{0}^{\pi}L^{j}\left(\kappa\cos\theta\right)\theta^{2}(\sin\theta)^{p-2}\mathrm{d}\theta\right$. (5.80)
$~~~~~~~\left.+\int_{\Omega_{v_{1}}}\int_{0}^{\pi}O(\theta^{3})L^{j}\left(\kappa\cos\theta\right)(\sin\theta)^{p-2}\mathrm{d}\theta\omega_{p-2}(\mathrm{d}\xi)\right\}\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{1})$ (5.81)
$=\psi_{4}(K)c_{j}(\kappa)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\mathrm{d}v+o(c_{j}(\kappa)),$ (5.82)
where the last line follows from a combination of (5.7) and (5.9). By similar arguments, we have
$h^{-1}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ $=\psi_{4}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f_{U}^{2}(u)\mathrm{d}u+o(1),$ (5.83)
which completes the proof. ∎
Appendix B Proof of main results
B.1 Proof of Proposition 2.1
Recalling (2.3) and the notation $h^{*}(u):=f_{U}(u)=\int_{\mathbb{S}^{p-1}}f(u,v)\ \omega_{p-1}(\mathrm{d}v)$, we obtain for any function $h:\mathbb{S}^{p-1}\to\mathbb{R}$
$\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h^{*}(u)f_{0}(v))(h^{*}(u)f_{0}(v)-h(u)f_{0}(v))\omega_{p-1}(\mathrm{d}v)$ (5.84)
$=(h^{*}(u)-h(u))\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h^{*}(u)f_{0}(v))\omega_{p-1}^{-1}\omega_{p-1}(\mathrm{d}v)$ (5.85)
$=\omega_{p-1}^{-1}(h^{*}(u)-h(u))(h^{*}(u)-h^{*}(u))=0.$ (5.86)
Then, if $hf_{0}$ is the density of a spherical symmetric distribution (for some density $h$ on $\mathbb{R}^{+}$), we obtain for the squared $L^{2}$-distance between $f$ and $hf_{0}$
$\Psi(h)$ $:=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.87)
$=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h^{*}(u)f_{0}(v)+h^{*}(u)f_{0}(v)-h(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.88)
$=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}\{(f(u,v)-h^{*}(u)f_{0}(v))^{2}+(h^{*}(u)f_{0}(v)-h(u)f_{0}(v))^{2}\}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.89)
$\geq\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h^{*}(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v).$ (5.90)
Therefore, $\Psi(h)$ is minimized for $h=h^{*}$, and
$\mathcal{M}^{2}$ $=\mathcal{M}^{2}(h^{*})$ (5.91)
$=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h^{*}(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.92)
$=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)-2\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{*}(u)f(u,v)f_{0}(v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.93)
$+\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f_{U}(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.94)
$=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)-\omega_{p-1}^{-1}\int_{\mathbb{R}^{+}}f_{U}^{2}(u)\mathrm{d}u.$ (5.95)
B.2 Martingale structure under spherical symmetry
Recall the representation of the statistic $\hat{\cal M}_{n}^{2}$ in (2.12), where the kernel $H_{n}$ is defined in (2.13), and let $\mathcal{F}_{i-1}=\sigma(Y_{1},\ldots,Y_{i-2},Y_{i-1})$ denote the sigma field generated by the random variables $Y_{1},\ldots,Y_{i-2},Y_{i-1}$. In the following, we show that in the case ${\cal M}^{2}=0$ the statistic $\hat{\cal M}_{n}$ is a cumulative sum of martingale differences with respect to the filtration $({\cal F}_{i})_{i=1}^{n}$. This is the cornerstone for the analysis of the asymptotic properties of the test statistic in the case ${\cal M}^{2}$=0.
Lemma B.1.
Suppose that Assumption 3.1 and 3.2 are satisfied and that $\mathcal{M}^{2}=0$, then the sequence
$\big(D_{i}\big)_{i=2}^{n}=\Big(\sum_{j=1}^{i-1}H_{n}(Y_{i},Y_{j})\Big)_{i=2}^{n}$ (5.96)
is a martingale difference sequence with respect tot the filtration $({\cal F}_{i}\big)_{i=1}^{n}$.
Proof.
By elementary calculation, we have (note that ${\cal M}^{2}=0$, which implies $f(u,v)=f_{U}(u)f_{0}(v)$)
$\mathbb{E}\big[D_{i}\big|{\cal F}_{i}\big]=\sum_{j=1}^{i-1}\mathbb{E}\big[H_{n}(Y_{i},Y_{j})\big|Y_{j}\big]$ (5.97)
$=h^{-1}\sum_{j=1}^{i-1}\left\{\int_{\mathbb{R}^{+}}K\left(\frac{U_{i}-u_{j}}{h}\right)f_{U}(u_{j})\mathrm{d}u_{j}\right\}\left\{\int_{\mathbb{S}^{p-1}}c_{1}^{-1}(\kappa)L(\kappa V_{i}^{\top}v_{j})f_{0}(v_{j})\ \omega_{p-1}(\mathrm{d}v_{j})-\omega_{p-1}^{-1}\right\}$ (5.98)
$=h^{-1}\omega_{p-1}^{-1}\sum_{j=1}^{i-1}\left\{\int_{\mathbb{R}^{+}}K\left(\frac{U_{i}-u_{j}}{h}\right)f_{U}(u_{j})\mathrm{d}u_{j}\right\}\left\{\int_{\mathbb{S}^{p-1}}c_{1}^{-1}(\kappa)L(\kappa V_{i}^{\top}v_{j})\ \omega_{p-1}(\mathrm{d}v_{j})-1\right\}$ (5.99)
Using the tangent-norm representation (3.2) with $y=v_{j}$, $x=V_{i}$ and (3.3) it follows that
$\mathbb{E}\big[D_{i}\big|{\cal F}_{i}\big]$ $=h^{-1}\omega_{p-1}^{-1}\sum_{j=1}^{i-1}\left\{\int_{\mathbb{R}^{+}}K\left(\frac{U_{i}-u_{j}}{h}\right)f_{U}(u_{j})\mathrm{d}u_{j}\right\}$ (5.100)
$\times\left\{\int_{\Omega_{V_{i}}}\int_{0}^{\pi}c_{1}^{-1}(\kappa)L(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-1\right\}$ (5.101)
$=h^{-1}\omega_{p-1}^{-1}\sum_{j=1}^{i-1}\left\{\int_{\mathbb{R}^{+}}K\left(\frac{U_{i}-u_{j}}{h}\right)f_{U}(u_{j})\mathrm{d}u_{j}\right\}$ (5.102)
$\times\left\{c_{1}^{-1}(\kappa)\omega_{p-2}\int_{0}^{\pi}L(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta-1\right\}=0.$ (5.103)
Since $D_{i}=\sum_{j=1}^{i-1}H_{n}(Y_{i},Y_{j})$ is $\mathcal{F}_{i}$ measurable, $(D_{i})_{i=2}^{n}$ is martingale difference sequence. ∎
B.3 Proof of Proposition 3.1
  • (i)
    If $\mathcal{M}^{2}=0$, the assertion follows from Lemma B.1.
  • (ii)
    If $\mathcal{M}^{2}>0$, we obtain from (2.12) and Lemma A.2
    $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})$ $=\frac{2}{n(n-1)}\sum_{i=1}^{n}\sum_{j=1}^{i-1}\mathbb{E}H_{n}(Y_{i},Y_{j})$ (5.104)
    $=\mathbb{E}(H_{n}(Y_{1},Y_{2}))$ (5.105)
    $=\frac{1}{hc_{1}(\kappa)}\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]-\frac{\omega_{p-1}^{-1}}{h}\mathbb{E}\Big[K\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ (5.106)
    $=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)-\omega_{p-1}^{-1}\int_{\mathbb{R}^{+}}f_{U}^{2}(u)\mathrm{d}u$ (5.107)
    $+\frac{h^{2}\phi_{2}(K)}{2}\left\{\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\left.\frac{\partial^{2}f(u,v_{1})}{\partial u^{2}}\right|_{u=u_{2}}\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{1})-\omega_{p-1}^{-1}\int_{\mathbb{R}^{+}}f_{U}(u)f_{U}^{\prime\prime}(u)\mathrm{d}u\right\}$ (5.108)
    $+\frac{\omega_{p-2}}{2\kappa}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f(u_{2},v_{1})\mathrm{tr}\{D^{2}_{f}(u_{2},v_{1})\}\mathrm{d}u_{2}\omega_{p-1}(\mathrm{d}v_{1})+O(\kappa^{-3/2}+h^{3}+h^{2}\kappa^{-1}).$ (5.109)
    and the result follows by an application of Lemma A.1.
B.4 Proof of Theorem 3.1
For the calculation of the asymptotic variance of $\hat{\mathcal{M}}^{2}_{n}$, note that for $j,k<i$ and $j\neq k$, we have
$\begin{split}\mathbb{E}(H_{n}(Y_{i},Y_{j})H_{n}(Y_{i},Y_{k}))&=\mathbb{E}\{H_{n}(Y_{i},Y_{j})\mathbb{E}(H_{n}(Y_{i},Y_{k})|(Y_{i},Y_{j}))\}\\ &=\mathbb{E}\{H_{n}(Y_{i},Y_{j})\mathbb{E}(H_{n}(Y_{i},Y_{k})|Y_{i})\}=0,\end{split}$
where the last equality follows by the same arguments as given in the proof of Lemma B.1. By the same Lemma $\hat{\cal M}_{n}^{2}$ is a sum of martingale differences, and therefore (B.4) implies
$s_{n}^{2}=\mathrm{Var}(\hat{\mathcal{M}}^{2}_{n})$ $=\frac{4}{n^{2}(n-1)^{2}}\mathrm{Var}\Big(\sum_{i=1}^{n}\sum_{j=1}^{i-1}H_{n}(Y_{i},Y_{j})\Big)$ (5.110)
$=\frac{4}{n^{2}(n-1)^{2}}\sum_{i=1}^{n}\mathrm{Var}\Big(\sum_{j=1}^{i-1}H_{n}(Y_{i},Y_{j})\Big)$ (5.111)
$=\frac{4}{n^{2}(n-1)^{2}}\sum_{i=1}^{n}\sum_{j=1}^{i-1}\mathrm{Var}(H_{n}(Y_{i},Y_{j}))$ (5.112)
$=\frac{2}{n(n-1)}\mathrm{Var}(H_{n}(Y_{i},Y_{j}))$ (5.113)
$=\frac{2}{n(n-1)h^{2}}\Big\{c_{1}^{-2}(\kappa)\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{2}(\kappa V_{i}^{\top}V_{j})\Big]$ (5.114)
$-2c^{-1}_{1}(\kappa)\omega_{p-1}^{-1}\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]+\omega_{p-1}^{-2}\mathbb{E}\Big[K^{2}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]\Big\}$ (5.115)
$=\frac{2}{n(n-1)h}\left[c_{1}^{-2}(\kappa)c_{2}(\kappa)\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)\right$. (5.116)
$\left.-2\omega_{p-1}^{-1}\psi_{2}(K)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)+\omega_{p-1}^{-2}\psi_{2}(K)\int_{\mathbb{R}^{+}}f_{U}^{2}(u)\mathrm{d}u\right$. (5.117)
$\left.+O(c_{1}^{-2}(\kappa)c_{2}(\kappa)h^{2}+c_{1}^{-1}(\kappa)b_{2}(\kappa))\right]$ (5.118)
$\sim\frac{2\psi_{2}(K)c_{1}^{-2}(\kappa)c_{2}(\kappa)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)}{n(n-1)h}$ (5.119)
$\sim\frac{2\psi_{2}(K)d^{-2}_{1}(p)d_{2}(p)\kappa^{(p-1)/2}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)}{n(n-1)h},$ (5.120)
where we have used Lemma A.2 three times in the last equality and Lemma A.1 for last two approximations.
By Lemma B.1, ${\cal M}_{n}^{2}$ is a sum of martingale difference and we can apply a central limit theorem for sums of martingale differences (Theorem 1 in Hall,, 1984) to prove the statement in Theorem 3.1. We have already shown that $H_{n}$ is symmetric, $\mathbb{E}\{H_{n}(Y_{1},Y_{2})|Y_{1}\}=0$ a.s., and by Proposition 3.1 it follows that $\mathbb{E}\{H_{n}^{2}(Y_{1},Y_{2})\}<\infty$ for each $n$. Therefore it is sufficient to verify that the condition
$[\mathbb{E}\{G_{n}^{2}(Y_{1},Y_{2})\}+n^{-1}\mathbb{E}\{H_{n}^{4}(Y_{1},Y_{2})\}]/[\mathbb{E}\{H_{n}^{2}(Y_{1},Y_{2})\}]^{2}\to 0$ (5.122)
is satisfied as $n\to\infty$, where
$G_{n}(x,y)=\mathbb{E}\big[H_{n}(Y_{1},x)H_{n}(Y_{1},y)\big].$ (5.123)
For this purpose note that
$h^{4}\mathbb{E}[H_{n}^{4}(Y_{1},Y_{2})]$ $=c_{1}^{-4}(\kappa)\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{4}(\kappa V_{i}^{\top}V_{j})\Big]-4c_{1}^{-1}(\kappa)\omega_{p-1}^{-3}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L(\kappa V_{i}^{\top}V_{j})\Big]$ (5.124)
$+6c_{1}^{-2}(\kappa)\omega_{p-1}^{-2}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{2}(\kappa V_{i}^{\top}V_{j})\Big]$ (5.125)
$-4c_{1}^{-3}(\kappa)\omega_{p-1}^{-1}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)L^{2}(\kappa V_{i}^{\top}V_{j})\Big]+\omega_{p-1}^{-4}\mathbb{E}\Big[K^{4}\Big(\frac{U_{i}-U_{j}}{h}\Big)\Big]$ (5.126)
and that
$\mathbb{E}\big[H_{n}^{2}(Y_{1},Y_{2})\big]={\rm Var}(H_{n}(Y_{1},Y_{2}))\sim\psi_{2}(K)c_{1}^{-2}(\kappa)c_{2}(\kappa)h^{-1}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}f^{2}(u,v)\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v),$ (5.127)
which follows from the calculation of $s_{n}^{2}$. Therefore, Lemma A.1 and A.3 yield
$n^{-1}\mathbb{E}\{H_{n}^{4}(Y_{1},Y_{2})\}/(\mathbb{E}(H_{n}^{2}(Y_{1},Y_{2}))^{2}=O(c_{4}(\kappa)/(c_{2}^{2}(\kappa)nh))=O(\kappa^{(p-1)/2}/(nh)).$ (5.128)
For the calculation of $G_{n}(x,y)$ we define $u_{2}=\|x\|$, $v_{2}=x/\|x\|$, $u_{3}=\|y\|$, $v_{3}=y/\|y\|$ and consider
$h^{2}G_{n}(x,y)$ $=h^{2}\mathbb{E}\big[H_{n}(Y_{1},x)H_{n}(Y_{1},y)\big]$ (5.129)
$=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{c_{1}^{-1}(\kappa)K\left(\frac{u_{1}-u_{2}}{h}\right)L(\kappa v_{1}^{\top}v_{2})-\omega_{p-1}^{-1}K\left(\frac{u_{1}-u_{2}}{h}\right)\right\}$ (5.130)
$\times\left\{c_{1}^{-1}(\kappa)K\left(\frac{u_{1}-u_{3}}{h}\right)L(\kappa v_{1}^{\top}v_{3})-\omega_{p-1}^{-1}K\left(\frac{u_{1}-u_{3}}{h}\right)\right\}f(u_{1},v_{1})\mathrm{d}u_{1}\ \omega_{p-1}(\mathrm{d}v_{1}).$ (5.131)
$=G_{1,n}(x,y)-G_{2,n}(x,y)-G_{3,n}(x,y)+G_{4,n}(x,y),$ (5.132)
where $G_{1,n}(x,y),\ldots,G_{4,n}(x,y)$ are defined by
$G_{1,n}(x,y)$ $=\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{c_{1}^{-2}(\kappa)K\left(\frac{u_{1}-u_{2}}{h}\right)L(\kappa v_{1}^{\top}v_{2})K\left(\frac{u_{1}-u_{3}}{h}\right)L(\kappa v_{1}^{\top}v_{3})\right\}f(u_{1},v_{1})\mathrm{d}u_{1}\ \omega_{p-1}(\mathrm{d}v_{1})$ (5.133)
$G_{2,n}(x,y)$ $=\omega_{p-1}^{-1}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{c_{1}^{-1}(\kappa)K\left(\frac{u_{1}-u_{2}}{h}\right)L(\kappa v_{1}^{\top}v_{2})K\left(\frac{u_{1}-u_{3}}{h}\right)\right\}\mathrm{d}u_{1}\ \omega_{p-1}(\mathrm{d}v_{1})$ (5.134)
$G_{3,n}(x,y)$ $=\omega_{p-1}^{-1}\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\left\{c_{1}^{-1}(\kappa)K\left(\frac{u_{1}-u_{2}}{h}\right)L(\kappa v_{1}^{\top}v_{3})K\left(\frac{u_{1}-u_{3}}{h}\right)\right\}\mathrm{d}u_{1}\ \omega_{p-1}(\mathrm{d}v_{1})$ (5.135)
$G_{4,n}(x,y)$ $=\omega_{p-1}^{-2}\int_{\mathbb{R}^{+}}K\left(\frac{u_{1}-u_{2}}{h}\right)K\left(\frac{u_{1}-u_{3}}{h}\right)f_{U}(u_{1})\mathrm{d}u_{1}.$ (5.136)
Note that we obtain from (5.11),
$L(\kappa t)\lesssim\kappa^{(p-1)/2}$ (5.137)
for all $t\in[-1,1]$, and with the representation
$v_{1}$ $=v_{2}\cos\theta+\xi\sin\theta,$ (5.138)
with $\xi\in\Omega_{v_{2}}:=\{\xi\in\mathbb{S}^{p-1}:~\xi\perp v_{2}\}$, it follows that
$|G_{1,n}(x,y)|\lesssim$ $\Big|h\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\Big\{c_{1}^{-2}(\kappa)K\left(u\right)L(v_{1}^{\top}v_{2})K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\kappa^{(p-1)/2}\Big\}$ (5.139)
$\times f(uh+u_{2},v_{1})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{1})\Big|$ (5.140)
$\lesssim$ $\Big|h\int_{\Omega_{v_{2}}}\int_{0}^{\pi}\int_{-1}^{1}\Big\{K\left(u\right)L(\cos\theta)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\kappa^{(p-1)/2}\Big\}$ (5.141)
$\times f(uh+u_{2},v_{2}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\Big|$ (5.142)
$\lesssim$ $h\kappa^{(p-1)/2}\int_{-1}^{1}\big|K(u)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\Big|du.$ (5.143)
We obtain for the terms $G_{n,2}$ and $G_{n,3}$ by similar arguments the estimate
$|G_{2,n}(x,y)+G_{3,n}(x,y)|$ $\leq h\omega_{p-1}^{-1}c_{1}^{-1}(\kappa)\int_{\mathbb{S}^{p-1}}\int_{\mathbb{R}^{+}}\Big|K\left(u\right)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\Big|\{L(\kappa v_{1}^{\top}v_{3})+L(\kappa v_{1}^{\top}v_{2})\}\mathrm{d}u\ \omega_{p-1}{\mathrm{d}v_{1}}$ (5.144)
$\lesssim h\kappa^{p/2-1/2}\int_{-1}^{1}\big|K(u)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\Big|du,$ (5.145)
and finally for the term $G_{4,n}$
$|G_{4,n}(x,y)|\leq h\omega_{p-1}^{-2}\int_{-1}^{1}\Big|K(u)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\Big|du.$ (5.146)
Then, combining (5.143), (5.145) and (5.146) and using Cauchy’s inequality yields
$h^{4}\mathbb{E}\{G_{n}^{2}(Y_{1},Y_{2})\}$ $\lesssim\mathbb{E}[G^{2}_{1,n}(Y_{1},Y_{2})]+\mathbb{E}\big[\{G_{2,n}(Y_{1},Y_{2})+G_{3,n}(Y_{1},Y_{2})\}^{2}\big]+\mathbb{E}[G^{2}_{4,n}(Y_{1},Y_{2})]$ (5.147)
$\lesssim h^{2}\kappa^{p-1}\int_{\mathbb{R}^{+}}\int_{\mathbb{R}^{+}}\Big\{\int_{-1}^{1}\Big|K\left(u\right)K\Big(u+\frac{u_{2}-u_{3}}{h}\Big)\Big|\mathrm{d}u\Big\}^{2}\mathrm{d}u_{2}\mathrm{d}u_{3}$ (5.148)
$\lesssim h^{3}\kappa^{p-1}\int_{-2}^{2}\Big\{\int_{-1}^{1}\big|K\left(u\right)K\left(u+a\right)\big|\mathrm{d}u\Big\}^{2}\mathrm{d}a$ (5.149)
$=O(h^{3}\kappa^{p-1}).$ (5.150)
Finally, with (5.127), we have
$\mathbb{E}\big[G_{n}^{2}(Y_{1},Y_{2})\big]\big/\big(\mathbb{E}\big[H_{n}^{2}(Y_{1},Y_{2}\big]\big)^{2}=O(h^{-1}\kappa^{p-1}/(h^{-2}\kappa^{p-1}))=O(h),$ (5.151)
and combining this statement with (5.128) yields (5.122), which completes the proof of Theorem 3.1.
B.5 Proof of Theorem 3.2
We recall the definition of the kernel
$H_{n}(Y_{i},Y_{j})=h^{-1}c_{1}^{-1}(\kappa)K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j})-h^{-1}\omega_{p-1}^{-1}K\left(\frac{U_{i}-U_{j}}{h}\right),$ (5.152)
introduce the notation
$g_{n}(Y_{1}):=\mathbb{E}\big[H_{n}(Y_{1},Y_{2})|Y_{1}\big]$ (5.153)
and define the random variables
$g(Y_{i})=f(U_{i},V_{i})-\omega_{p-1}^{-1}f_{U}(U_{i}),\ \ (i=1,\ldots,n)$ (5.154)
At the end of the proof, we show that
$R_{n}=\hat{\mathcal{M}}_{n}^{2}-\mathbb{E}[\hat{\mathcal{M}}_{n}^{2}]-\frac{2}{n}\sum_{i=1}^{n}\big(g(Y_{i})-\mathcal{M}^{2}\big)=o_{\mathbb{P}}\Big({1\over\sqrt{n}}\Big).$ (5.155)
Therefore, we shall first focus on deriving the limiting behavior of $\frac{1}{n}\sum_{i=1}^{n}(g(Y_{i})-\mathcal{M}^{2})$. By elementary calculation, we obtain
$\mathbb{E}[g(Y_{i})]=\mathcal{M}^{2}~,~~\mathrm{Var}(g(Y_{i}))=\sigma^{2},$ (5.156)
where ${\cal M}^{2}$ and $\sigma^{2}$ are defined in (2.4) and (3.12) respectively. Therefore, since $Y_{1},\ldots,Y_{n}$ are independent identically distributed, the Central Limit Theorem yields
$\sqrt{n}\Big\{\frac{1}{n}\sum_{i=1}^{n}g(Y_{i})-\mathcal{M}^{2}\Big\}=\sqrt{n}\Big\{\frac{1}{n}\sum_{i=1}^{n}\big(g(Y_{i})-\mathbb{E}[g(Y_{i})]\big)\Big\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,\sigma^{2}).$ (5.157)
Finally, we prove (5.155). For this purpose, note that
$R_{n}$ $=\hat{\mathcal{M}}_{n}^{2}-\frac{2}{n}\sum_{i=1}^{n}g(Y_{i})+2\mathcal{M}^{2}-\mathbb{E}[H_{n}(Y_{1},Y_{2})]$ (5.158)
$=r_{1,n}+2r_{2,n},$ (5.159)
where
$r_{1,n}$ $=\frac{1}{n(n-1)}\sum_{i\neq j=1}^{n}\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}(H_{n}(Y_{i},Y_{j}))\big\}$ (5.160)
$r_{2,n}$ $=\frac{1}{n}\sum_{i=1}^{n}\{(g_{n}(Y_{i})-\mathbb{E}(H_{n}(Y_{i},Y_{j})))-(g(Y_{i})-\mathcal{M}^{2})\}.$ (5.161)
By triangle inequality, we have
$\|R_{n}\|:=\big\{\mathbb{E}[R^{2}_{n}]\big\}^{1/2}\leq\|r_{1,n}\|+2\|r_{2,n}\|.$ (5.162)
and we show at the end of the proof that
$\|r_{j,n}\|=o\big(n^{-1/2}\big)~~~~(j=1,2).$ (5.163)
Then $\|R_{n}\|=o(n^{-1/2})$ and, by Markov’s inequality, $R_{n}=o_{\mathbb{P}}\big({1\over\sqrt{n}}\big)$, which proves (5.155). Combining this result with (5.157) yields the assertion of of Theorem 3.2, that is
$\sqrt{n}(\hat{\mathcal{M}}_{n}^{2}-\mathbb{E}[\hat{\mathcal{M}}_{n}^{2}])\Rightarrow N(0,4\sigma^{2}).$ (5.164)
The proof is now completed proving the estimates in (5.163). To derive a corresponding estimate for $r_{1,n}$ we note that, by conditioning on $Y_{i},Y_{j}$,
$\mathrm{Cov}\big(\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}(H_{n}[(Y_{i},Y_{j})]\big\},\big\{H_{n}(Y_{i},Y_{k})-g_{n}(Y_{i})-g_{n}(Y_{k})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\}]=0,$
whenever the indices $i,j,k$ are different from each other. Consequently, we obtain from Proposition 3.1
$\mathrm{Var}(r_{1,n})$ $=\frac{1}{n(n-1)}\mathrm{Var}\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\}$ (5.165)
$+\frac{n-2}{n(n-1)}\mathrm{Cov}[\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\},$ (5.166)
$\{H_{n}(Y_{i},Y_{k})-g_{n}(Y_{i})-g_{n}(Y_{k})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\}]$ (5.167)
$=\frac{1}{n(n-1)}\big[\mathrm{Var}\{H_{n}(Y_{i},Y_{j})\}-2\mathrm{Var}\{g_{n}(Y_{i})\}\big]$ (5.168)
$=O\Big(s_{n}^{2}+\frac{1}{n(n-1)}\sigma^{2}\Big)=O(\kappa^{(p-1)/2}n^{-2}h^{-1}+n^{-2})=o(n^{-1}),$ (5.169)
where the second equality follows by a tedious calculation observing that that
$\mathbb{E}\{H_{n}(Y_{i},Y_{j})H_{n}(Y_{i},Y_{k})\}$ $=\mathbb{E}\{H_{n}(Y_{i},Y_{j})\mathbb{E}(H_{n}(Y_{i},Y_{k})|(Y_{i},Y_{j}))\}$ (5.170)
$=\mathbb{E}\{H_{n}(Y_{i},Y_{j})\mathbb{E}(H_{n}(Y_{i},Y_{k})|Y_{i})\}$ (5.171)
$=\mathbb{E}\{H_{n}(Y_{i},Y_{j})g_{n}(Y_{i})\}$ (5.172)
$=\mathbb{E}(\mathbb{E}(H_{n}(Y_{i},Y_{j})|Y_{i})\mathbb{E}\{H_{n}(Y_{i},Y_{k})|Y_{i})\}=\mathbb{E}(g_{n}^{2}(Y_{i})),$ (5.173)
whenever the indices $i,j,k$ are different. As $r_{1,n}$ is centered, (5.163) follows for $j=1$. Finally, since $Y_{i}$’s are $i.i.d.$, we have
$\|r_{2,n}\|\lesssim n^{-1/2}\|(g_{n}(Y_{i})-\mathbb{E}(H_{n}(Y_{i},Y_{j})))-(g(Y_{i})-\mathcal{M}^{2})\|=O(n^{-1/2}(1/\kappa+h^{2}))=o(n^{-1/2}).$ (5.174)
B.6 Proof of Theorem 4.1
Recall the definition of the process $S_{n}$ in (4.1), define the random variables
$g(Y_{i})=f(U_{i},V_{i})-\omega_{p-1}^{-1}f_{U}(U_{i})$ (5.175)
($i=1,\ldots,n$) and consider the stochastic process
$S_{n}^{\circ}(t)=\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}(g(Y_{i})-\mathcal{M}^{2}).$ (5.176)
The assertion of Theorem 4.1 is now proved in two steps.
Step 1: we show that
$\max_{{2}\leq k\leq n}{1\over\sqrt{n}}\Big|k(\hat{\mathcal{M}}_{k}^{2}-\mathbb{E}[\hat{\mathcal{M}}_{n}^{2}])-2\sum_{i=1}^{k}(g(Y_{i})-\mathcal{M}^{2})\Big|=o_{\mathbb{P}}(1),$ (5.177)
which implies
$\sup_{t\in[0,1]}|S_{n}(t)-S_{n}^{\circ}(t)|=o_{\mathbb{P}}(1).$ (5.178)
Step 2: we prove that
$\big\{S_{n}^{\circ}(t)\big\}_{t\in[0,1]}\Rightarrow\big\{\sigma\mathbb{B}(t)\big\}_{t\in[0,1]}$ (5.179)
in $\ell^{\infty}([0,1])$, where $\big\{\mathbb{B}(t)\big\}_{t\in[0,1]}$ is a standard Brownian motion and $\sigma^{2}$ is defined in (3.12).
Proof of Step 1: Recall that
$g_{n}(Y_{1})=\mathbb{E}[H_{n}(Y_{1},Y_{2})|Y_{1})],$ (5.180)
and note that
$R_{k,n}^{\circ}$ $=k(\hat{\mathcal{M}}_{k}^{2}-\mathbb{E}[\hat{\mathcal{M}}_{n}^{2}])-2\sum_{i=1}^{k}(g(Y_{i})-\mathcal{M}^{2})$ (5.181)
$=\frac{1}{(k-1)}\sum_{i\neq j=1}^{k}\Big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]$ (5.182)
$+g_{n}(Y_{i})-\mathbb{E}[H_{n}(Y_{i},Y_{j})]+g_{n}(Y_{j})-\mathbb{E}[H_{n}(Y_{i},Y_{j})]-(g(Y_{i})-\mathcal{M}^{2})-(g(Y_{j})-\mathcal{M}^{2})\Big\}.$
For a real valued random variable $Z$ we define $\|Z\|$ denote $\{\mathbb{E}[Z^{2}]\}^{1/2}$, then we obtain
$\Big\|\max_{2\leq k\leq n}|R_{k,n}^{\circ}|\Big\|$ $\leq R_{1}+2R_{2},$ (5.183)
where $R_{1}$ and $R_{2}$ are defined by
$R_{1}$ $=\Big\|\max_{2\leq k\leq n}\Big|\frac{1}{(k-1)}\sum_{i\neq j=1}^{k}\Big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\Big\}\Big|\Big\|$ (5.184)
$R_{2}$ $=\Big\|\max_{2\leq k\leq n}\Big|\sum_{i=1}^{k}\big\{g_{n}(Y_{i})-\mathbb{E}[H_{n}(Y_{i},Y_{j})]-(g(Y_{i})-\mathcal{M}^{2})\big\}\Big|\Big\|.$ (5.185)
Similar arguments as given in the proof of Lemma B.1 show that the random variables
$\sum_{j=1}^{i-1}\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\}$
are martingale differences with respect to the filtration $(\mathcal{F}_{i})_{i=1,\ldots,n}$, where $\mathcal{F}_{i}=\sigma(Y_{1}\ldots,Y_{i})$ is the sigma field generated by $Y_{1}\ldots,Y_{i}$. Moreover, using similar arguments as in the calculation in (5.169), we obtain
$\mathrm{Var}\Big(\sum_{j=1}^{i-1}\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\}\Big)$ (5.186)
$=\sum_{j=1}^{i-1}\mathrm{Var}\big(\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big)$ (5.187)
$+\sum_{j\neq j^{\prime}}^{i-1}\mathrm{Cov}\Big(\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\},$ (5.188)
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\big\{H_{n}(Y_{i},Y_{j^{\prime}})-g_{n}(Y_{i})-g_{n}(Y_{j^{\prime}})+\mathbb{E}[H_{n}(Y_{i},Y_{j^{\prime}})]\big\}\Big)$ (5.189)
$=(i-1)[\mathrm{Var}\{H_{n}(Y_{1},Y_{2})\}-2\mathrm{Var}\{g_{n}(Y_{1})\}]$ (5.190)
$=O((i-1)(\kappa^{(p-1)/2}h^{-1}+\sigma^{2})).$ (5.191)
For a constant $\rho>0$, $c=\lfloor 1/(\log\rho)\rfloor+1$, we have
$R_{1}$ $\leq\Big\|\max_{1\leq l\leq\lfloor c\log n\rfloor}\max_{\rho^{l-1}\leq k\leq\rho^{l}}\Big|\frac{1}{(k-1)}\sum_{i\neq j=1}^{k}\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\}\Big|\Big\|$ (5.192)
$\leq\sum_{l=1}^{\lfloor c\log n\rfloor}(\rho^{l-1}-1)^{-1}\Big\|\max_{1\leq k\leq\rho^{l}}\Big|\sum_{i\neq j=1}^{k}\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\}\Big|\Big\|$ (5.193)
$=O\Big(\sum_{l=1}^{\lfloor c\log n\rfloor}\rho^{-l+1}\Big\|\sum_{i=1}^{\rho^{l}}\sum_{j=1}^{i-1}\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\}\Big\|\Big)$ (5.194)
$=O\Big(\sum_{l=1}^{\lfloor c\log n\rfloor}\rho^{-l+1}\Big\{\sum_{i=1}^{\rho^{l}}\Big\|\sum_{j=1}^{i-1}\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\}\|^{2}\Big\}^{1/2}\Big)$ (5.195)
$=O\Big(\sum_{l=1}^{\lfloor c\log n\rfloor}\rho^{-l+1}\Big\{\sum_{i=1}^{\rho^{l}}(i-1)(\kappa^{(p-1)/2}h^{-1}+\sigma^{2})\Big\}^{1/2}\Big)$ (5.196)
$=O\big(\log n(\kappa^{(p-1)/2}h^{-1})^{1/2})\big)=o(\sqrt{n}),$ (5.197)
where the first equality follows from Doob’s inequality, the second follows from the fact that $\sum_{j=1}^{i-1}\big\{H_{n}(Y_{i},Y_{j})-g_{n}(Y_{i})-g_{n}(Y_{j})+\mathbb{E}[H_{n}(Y_{i},Y_{j})]\big\}$ are martingale differences and the third follows from (5.191). Similarly, applying Doob’s inequality for the $i.i.d.$ sequence $g_{n}(Y_{i})-g(Y_{i})$
$R_{2}$ $\leq\Big\|\sum_{i=1}^{n}\{g_{n}(Y_{i})-\mathbb{E}[H_{n}(Y_{i},Y_{j})]-(g(Y_{i})-\mathcal{M}^{2})\}\Big\|$ (5.198)
$\lesssim\sqrt{n}\|g_{n}(Y_{i})-\mathbb{E}[H_{n}(Y_{i},Y_{j})]-(g(Y_{i})-\mathcal{M}^{2})\|\big)$ (5.199)
$\lesssim\sqrt{n}R_{21}+\sqrt{n}R_{22}+\sqrt{n}R_{23},$ (5.200)
where
$R_{21}$ $=\Big\|\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{-1}c_{1}^{-1}(\kappa)K\Big(\frac{u_{2}-U_{1}}{h}\Big)L(\kappa v_{2}^{\top}V_{1})f(u_{2},v_{2})\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|,$ (5.201)
$R_{22}$ $=\Big\|\omega_{p-1}^{-1}\Big\{\int_{\mathbb{R}^{+}}h^{-1}K\Big(\frac{u_{2}-U_{1}}{h}\Big)f_{U}(u_{2})\mathrm{d}u_{2}-f_{U}(U_{1})\Big\}\Big\|~,$ (5.202)
$R_{23}$ $=|\mathbb{E}[H_{n}(Y_{i},Y_{j})]-\mathcal{M}^{2}|.$ (5.203)
A Taylor expansion gives for the first term
$R_{21}$ $=\Big\|\int_{-1}^{1}\int_{\mathcal{\mathbb{}}{S}^{p-1}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa v_{2}^{\top}V_{1})f(U_{1}+hu,v_{2})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|$ (5.204)
$\leq\Big\|\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{v_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)$ (5.205)
$~~~~~~~~~~~~~~~~~~\times f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-f_{U,V}(U_{1},V_{1})\Big\|$ (5.206)
$\lesssim\Big\|\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{V_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)\big\{f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)$ (5.207)
$~~~~~~~~~~~~~~~~~~-f(U_{1},V_{1}\cos\theta+\xi\sin\theta)\big\}(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\Big\|$ (5.208)
$+\Big\|\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{V_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)\big\{f(U_{1},V_{1}\cos\theta+\xi\sin\theta)$ (5.209)
$~~~~~~~~~~~~~~~~~~-f(U_{1},V_{1})\big\}(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)\Big\|$ (5.210)
$\lesssim h+\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{V_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)\theta(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)$ (5.211)
$=O\big(h+b_{1}(\kappa)/c_{1}(\kappa)\big)=O(h+\kappa^{-1/2}),$ (5.212)
where we used the fact that $\partial f(u,v)/\partial u$ and $\mathcal{D}_{f}(u,v)$ are uniformly bounded and Lemma A.1. By similar but simpler arguments we obtain $R_{22}=O(h)$. By Proposition 3.1, we have, $R_{23}=O(1/\kappa+h^{2})$. Combining these estimates with (5.200) yields
$R_{2}=O(\sqrt{n}(h+\kappa^{-1/2}))=o(\sqrt{n}).$ (5.213)
Finally, (5.177) follows from (5.181), (5.197), (5.213).
Proof of Step 2: Note that the random variables $g(Y_{i})$ are $i.i.d.$ with
$\mathbb{E}[g(Y_{1})]=\mathcal{M}^{2}~~~\text{ and }~~\mathrm{Var}(g(Y_{1}))=\sigma^{2}.$ (5.214)
Therefore, by Donsker’s Theorem (see the discussion on page 225 - 226 in Van Der Vaart and Wellner,, 1996) it follows that
$\big\{S_{n}^{\circ}(t)\big\}_{t\in[0,1]}\Rightarrow\big\{\sigma\mathbb{B}(t)\big\}_{t\in[0,1]}$ (5.215)
in $\ell^{\infty}([0,1])$.

### v2

@@
 Abstract
-Most of the work on checking spherical symmetry assumptions on the distribution of $p$-dimensional random vector $Y$ has its focus on statistical tests for the null hypothesis of exact spherical symmetry.
-In this paper we take a different point of view and propose a measure for the deviation from spherical symmetry, which is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a random vector $Y_{s}$ with a spherical distribution.
-We develop estimators for the minimum distance with corresponding statistical guarantees (provided by asymptotic theory) and demonstrate the applicability of our approach by means of a simulation study and a real data example.
+Most of the work on checking spherical symmetry assumptions on the distribution of a $p$-dimensional random vector $Y$ focuses on testing for the null hypothesis of exact spherical symmetry.
+In this paper, we take a different point of view and consider the problem of testing if the deviation from spherical symmetry exceeds a given threshold.
+For this purpose, we propose a novel measure for the deviation from spherical symmetry, which is based on the minimum distance between the distribution of the vector $\big(|Y|,Y/|Y|)^{\top}$ and its best approximation by a distribution of a vector $\big(|Y_{s}|,Y_{s}/|Y_{s}|)^{\top}$ , where $Y_{s}$ follows a spherical distribution.
+We develop estimators for the minimum distance with corresponding statistical guarantees (provided by asymptotic theory) and use these results to develop statistical inference.
+Our approach assumes neither Gaussianity nor the existence of moments, and we demonstrate the applicability of our approach by means of a simulation study and a real data example.
 1 Introduction
-Spherical contoured models include multivariate Gaussian distribution, multivariate $t$- distribution and multivariate stable distribution find their applications in earth science, vector cardiology (Mardia and Jupp,, 1999; Di Marzio et al.,, 2019), portfolio theory (Gupta et al.,, 2013), Mahalanobis distances (Luis Angel García-Escudero,, 2005).
-There has been a wide interest in testing for the spherical symmetry, or a wider class of elliptical symmetry, see for example Smith, (1977), Baringhaus, (1991), Fang et al., (1993), Koltchinskii and Li, (1998), Schott, (2002), Huffer and Park, (2007), Liang et al., (2008), Cassart et al., (2008), Henze et al., (2014), Albisetti et al., (2020), Babić et al., (2021), Banerjee and Ghosh, (2024) among others.
+Spherical contoured models include multivariate Gaussian distribution, multivariate $t$- distribution and multivariate stable distribution find their applications in vector cardiology (mardia1999directional), portfolio theory (gupta2013elliptically), Mahalanobis distances (luis2005).
+There has been a wide interest in testing for the spherical symmetry, or a wider class of elliptical symmetry, see for example Smith01011977, ludwig1991, FANG199334, KOLTCHINSKII1998228, SCHOTT2002, HP2007, liangetal2008, PG2008, Henzeetal2014, holzman2020, babic2021, huang2023multivariate, banerjee2024consistent among others.
 A common feature of most of the cited references is that the authors use different approaches to construct tests for the null hypothesis of exact spherical symmetry.
-However, this assumption is rarely met because one believes that it is exactly (with mathematical equality) satisfied, but with the hope that the deviation from the null hypothesis is small such that (optimal) inference under the assumption of spherical symmetry is still reliable.
+However, this assumption is rarely met because one does not believe that it is exactly (with mathematical equality) satisfied, but with the hope that the deviation from the null hypothesis is small such that (optimal) inference under the assumption of spherical symmetry is still reliable.
 Thus, strictly speaking, there are many situations where tests for exact sphericity are performed, although it is clear that this hypothesis is at most “approximately” satisfied.
 In this paper, we take a different point of view on the problem of validating the assumption of spherical symmetry on the distribution of a random vector $Y$.
-Instead of testing for spherical symmetry, we propose a measure of deviation from spherical symmetry.
+Instead of testing for exact spherical symmetry, we propose a measure of deviation from spherical symmetry.
 This measure is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a vector $Y_{s}$ with a spherical distribution, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{p}$ (note that a random variable $Y_{s}$ has a spherical distribution if and only if $\|Y_{s}\|$ and $Y_{s}/\|Y_{s}\|$ are independent and $Y_{s}/\|Y_{s}\|$ is uniformly distributed on the $p$-dimensional sphere $\mathbb{S}_{p-1}$).
@@
 Consequently, the estimation problem of the minimum distance boils down to estimating the squared $L^{2}$-norm of the density of a distribution on the sphere and a distribution on $\mathbb{R}^{+}$.
-While the problem of estimating the integrated squared density on $\mathbb{R}^{p}$ has been extensively studied from various perspectives including optimality and adaptivity (see Hall and Marron,, 1987; Bickel and Ritov,, 1988; Laurent and Massart,, 2000; Giné and Nickl,, 2008, among many others), the corresponding problem on the sphere have not found much attention so far.
+While the problem of estimating the integrated squared density on $\mathbb{R}^{p}$ has been extensively studied from various perspectives including optimality and adaptivity (see hall1987estimation; bickel1988estimating; LaurantMassart2000; GineNickl2008, among many others), the corresponding problem on the sphere have not found much attention so far.
 We construct an estimator for the minimum distance from a sequence of independent identically distributed observations combining an estimator for the integrated squared density of the joint distribution of $Y/\|Y\|$ and $\|Y\|$ with an estimator for the integrated squared density of $\|Y\|$ on the non-negative line.
 We prove asymptotic normality of an appropriately centered and normalized version.
-Interestingly, from the theoretical perspective, there appear to be different scalings, depending on whether the distribution is spherical symmetric or not.
-In the latter case, the difference of the estimators converges at a rate $\sqrt{n}$ as in the case estimating the integral of the squared density (see, for example, Bickel and Ritov,, 1988).
-On the other hand, if the distribution of $Y$ is in fact spherical symmetric, the estimator converges at a rate $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, where $h\to 0$ and $\kappa\to\infty$ denote smoothing parameters required for the estimation of the integrated squared densities of $\|Y\|$ and $Y/\|Y\|$, respectively.
+Interestingly, from the theoretical perspective, there appear to be different scalings, depending on whether the distribution is spherically symmetric or not.
+In the latter case, the difference of the estimators converges at a rate $\sqrt{n}$ as in the case estimating the integral of the squared density (see, for example, bickel1988estimating).
+On the other hand, if the distribution of $Y$ is in fact spherically symmetric, the estimator converges at a rate $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, where $h\to 0$ and $\kappa\to\infty$ denote smoothing parameters required for the estimation of the integrated squared densities of $\|Y\|$ and $Y/\|Y\|$, respectively.
 Based on these results, we develop several statistical applications.
@@
 Moreover, as an alternative, we also develop in Section 4 pivotal inference for the measure ${\cal M}^{2}$, which does not require any estimation of asymptotic variances.
-Our approach is based on the self-normalization principle (see Shao,, 2015), which cannot be used directly in the present context, but has to be adapted to the specific statistical problems under consideration.
-In particular, we have to prove the weak convergence of a sequential estimator of ${\cal M}^{2}$ to obtain a pivotal statistic which can be used for the construction of confidence intervals and for testing the hypotheses of a relevant deviation from sphericity.
+Our approach is based on the self-normalization principle (see shao2015), which cannot be used directly in the present context.
+In particular, we have to derive the weak convergence of a process of non-parametric U-statistics with random elements on the sphere, and adapt it to obtain a pivotal statistic for the estimator of ${\cal M}^{2}$.
 Finally, we demonstrate the applicability of our approach through a small simulation study and a real data example.
@@
 for some function $g$.
-Consequently, observing (2.1), it follows that $Y$ has a spherical symmetric distribution if and only $U=\|Y\|$ and $V=Y/\|Y\|$ are independent and $V=Y/\|Y\|$ is uniformly distributed on the sphere with (constant) density
+Consequently, observing (2.1), it follows that $Y$ has a spherically symmetric distribution if and only if $U=\|Y\|$ and $V=Y/\|Y\|$ are independent and $V=Y/\|Y\|$ is uniformly distributed on the sphere with (constant) density
 $f_{0}(v)=\omega_{p-1}^{-1}I(\|v\|=1),$ (2.3)
@@
 By Proposition 2.1 the problem of estimating the minimum distance ${\cal M}^{2}$ boils down to the estimation of the integrated squared densities of the random variables $(U_{i},V_{i})$ and $U_{i}$.
-The latter can be estimated by standard methods (see, for example Hall and Marron,, 1987; Bickel and Ritov,, 1988, among many others) and we use
+The latter can be estimated by standard methods (see, for example hall1987estimation; bickel1988estimating, among many others) and we use
 $\hat{\cal M}_{n}^{(1)}:=\frac{1}{n(n-1)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)$ (2.6)
-as estimate for the integrated squared density of $U$, where $h$ denotes a bandwidth.
+as estimate for the integrated squared density of $U$, where $h$ denotes a bandwidth, $K$ is a symmetric kernel function supported on (1,1)(-1,1) with integral $1$, for example the Epanechnikov kernel.
 The estimation of the first integral in (2.5) is more intricate as it corresponds to the (squared) $L^{2}$-norm of a density on $\mathbb{R}^{+}\times\mathbb{S}^{p-1}$.
-For this purpose, we use classical results from density estimation on the sphere, such as Hall et al., (1987), and consider a rapidly varying function $L:\mathbb{R}^{p}\to\mathbb{R}$, which satisfies (i) $L$ is nonnegative and nondecreasing, (ii) for each $0<r<1$,
+For this purpose, we use classical results from density estimation on the sphere, such as HallWatsonCabrera1987, and consider a rapidly varying function $L:\mathbb{R}^{p}\to\mathbb{R}$, which satisfies (i) $L$ is nonnegative and nondecreasing, (ii) for each $0<r<1$,
 $L(rt)/L(t)\to 0,\ \text{as}\ t\to\infty.$ (2.7)
 The estimator of the first integral in Proposition 2.1 is then defined by
-$\hat{\cal M}_{n}^{(2)}:=\frac{1}{n(n-1)c_{1}(\kappa)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j})$ (2.8)
-where $K$ is a kernel function and $h$ is the same bandwidth as used in the estimator $\hat{\cal M}_{n}^{(1)}$, $\kappa\to\infty$ is another bandwidth corresponding to the spherical part of the vector $(U,V)$ and the constant $c_{1}(\kappa)$ is given by
+$\hat{\cal M}_{n}^{(2)}:=\frac{1}{n(n-1)c_{1}(\kappa)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j}),$ (2.8)
+where $K(\cdot)$ and $h$ are the same kernel and bandwidth as used in $\hat{\cal M}_{n}^{(1)}$, $\kappa\to\infty$ is another bandwidth corresponding to the spherical part of the vector $(U,V)$, and the constant $c_{1}(\kappa)$ is given by
 $c_{1}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta.$ (2.9)
@@
 $L(\kappa t)=\kappa^{p/2-1}\{(2\pi)^{p/2}\mathcal{I}_{p/2-1}(\kappa)\}^{-1}e^{\kappa t}$ (2.10)
-for the kernel $L$, where $\mathcal{I}_{\nu}$ is the modified Bessel function of the first kind (see also 9.6.18 on Page 376 of Abramowitz and Stegun,, 1968) of order $\nu$, i.e.,
+for the kernel function $L(\cdot)$, where $\mathcal{I}_{\nu}$ is the modified Bessel function of the first kind (see also 9.6.18 on Page 376 of abramowitz1968handbook) of order $\nu$, i.e.,
 $\mathcal{I}_{\nu}(\kappa)=\frac{(\kappa/2)^{\nu}}{\Gamma(\nu+1/2)\Gamma(1/2)}\int_{-1}^{1}e^{\kappa t}(1-t^{2})^{\nu-1/2}\mathrm{d}t.$ (2.11)
@@
 3.1 Asymptotic theory
-For the investigation of the bias of $\hat{\mathcal{M}}_{n}^{2}$, we recall a Taylor expansion for a twice differentiable function $f:\mathbb{R}^{+}\times\mathbb{S}^{p-1}\to\mathbb{R}$ with respect to the argument $v$ on the sphere $\mathbb{S}^{p-1}$.
-According to equation (3) in Marco Di Marzio and Taylor, (2014), we have
-$f(u,y)=f(u,x)+\theta\xi^{\top}\mathcal{D}_{f}(u,x)+\frac{\theta^{2}}{2}\xi^{\top}\mathcal{D}^{2}_{f}(u,x)\xi+O(\theta^{3}).$ (3.1)
-Here $D_{f}^{s}(u,x)$ is the matrix of the $s$th-order derivative of the function $f(u,v)$ with respect to $v$ at the point $x$ and for $x,y\in\mathbb{S}^{p-1}$, the vector $\xi$ and the angle $\theta\in(0,\pi)$ are defined by the tangent–normal decomposition
+For the investigation of the bias of $\hat{\mathcal{M}}_{n}^{2}$, we recall a Taylor expansion for a twice continuously differentiable function $f:\mathbb{R}^{+}\times\mathbb{S}^{p-1}\to\mathbb{R}$ with respect to the argument $v$ on the sphere $\mathbb{S}^{p-1}$.
+According to equation (3) in DiMarzio03042014, we have
+$f(u,y)=f(u,x)+\theta\xi^{\top}\mathcal{D}_{f}(u,x)+\frac{\theta^{2}}{2}\xi^{\top}\mathcal{D}^{2}_{f}(u,x)\xi+O(\theta^{3}),$ (3.1)
+where $\mathcal{D}_{f}^{s}(u,x)$ is the matrix of the $s$th-order derivative of the function $f(u,v)$ with respect to $v$ at the point $x$ and for $x,y\in\mathbb{S}^{p-1}$, the vector $\xi$ and the angle $\theta\in(0,\pi)$ are defined by the tangent–normal decomposition
 $y=x\cos(\theta)+\xi\sin(\theta).$ (3.2)
-Note that $\theta\in(0,\pi)$ is the angle between $x$ and $y$, and $\xi\in\mathbb{S}^{p-1}$ is a vector orthogonal to $x$, i.e., $\xi\in\Omega_{x}=\{\xi\in\mathbb{S}^{p-1}:\xi\perp x\}$.
+Note that $\theta\in(0,\pi)$ is the angle between $x$ and $y$, and $\xi\in\mathbb{S}^{p-1}$ is a vector orthogonal to $x$, i.e., $\xi\in\Omega_{x}=\{z\in\mathbb{S}^{p-1}:z\perp x\}$.
 The intuition for (3.1) comes from the fact that $y-x\approx\theta\xi$ when $\theta$ approximates 0, which follows from a Taylor expansion of the function $\cos\theta$ and $\sin\theta$ at the point $\theta=0$.
@@
 $\omega_{p-1}(\mathrm{d}y)=(\sin\theta)^{p-2}\mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi),$ (3.3)
-see equation (2) in Marco Di Marzio and Taylor, (2014) or equation (1.5) in Hall et al., (1987). On the other hand, for a fixed $v\in\mathbb{S}^{p-1}$, we have the Taylor expansion for $f$
+see equation (2) in DiMarzio03042014 or equation (1.5) in HallWatsonCabrera1987. Meanwhile, for a fixed $v\in\mathbb{S}^{p-1}$, we have the Taylor expansion for $f$
 $f(u_{0}+h,v)=f(u_{0},v)+h\left.\frac{\partial f(u,v)}{\partial u}\right|_{u=u_{0}}+\frac{h^{2}}{2}\left.\frac{\partial^{2}f(u,v)}{\partial u^{2}}\right|_{u=u_{0}}+O(h^{3}),$ (3.4)
-where $\frac{\partial^{s}f(u,v)}{\partial u^{s}}$ denotes the partial derivative of order $s$ with respect to $u$.
-Finally, we denote first and second derivative of the marginal density $f_{U}$ by $f_{U}^{\prime}$ and $f_{U}^{\prime\prime}$, respectively.
+where $\partial^{s}f(u,v)/\partial u^{s}$ denotes the partial derivative of order $s$ with respect to $u$.
+Finally, we denote first and second derivatives of the marginal density $f_{U}$ by $f_{U}^{\prime}$ and $f_{U}^{\prime\prime}$, respectively.
 For a precise statement of our results we require the following assumptions.
 Assumption 3.1.
-$K(\cdot)$ is a symmetric density function supported on the interval (1,1)(-1,1), and $\int_{-1}^{1}K^{4}\left(u\right)\mathrm{d}u<\infty$; the kernel $L(\cdot)$ is the Langevin kernel defined in (2.10).
+$K(\cdot)$ is a symmetric function supported on the interval (1,1)(-1,1), $\int_{-1}^{1}K\left(u\right)=1$ and $\int_{-1}^{1}K^{4}\left(u\right)\mathrm{d}u<\infty$; the kernel $L(\cdot)$ is the Langevin kernel defined in (2.10).
 Assumption 3.2.
-The density $f$ of $Y$ is twice differentiable.
+The density $f$ of $Y$ is twice continuously differentiable.
   • (a)
@@
   • (i)
-    If the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
+    If the distribution of $Y$ is spherically symmetric, that is $\mathcal{M}^{2}=0$, we have
     $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})=0.$ (3.5)
   • (ii)
-    If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
+    If the distribution of $Y$ is not spherically symmetric, that is $\mathcal{M}^{2}>0$, we have
     $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})$ $={\mathcal{M}}^{2}$ (3.6)
@@
     $\phi_{j}(K)=\int_{-1}^{1}x^{j}K(u)du.$ (3.9)
-By Proposition 3.1, the statistic $\hat{\mathcal{M}}^{2}_{n}$ is in general not an unbiased estimator of $\mathcal{M}^{2}$.
-The bias of the kernel estimation with respect to the radius $\|Y\|$ is of order $h^{2}$ and it is of order $1/\kappa$ with respect to the direction $Y/\|Y\|$.
-Therefore, smaller values of $h$ and larger values of $\kappa$ are preferred to diminish the bias the estimator $\hat{\mathcal{M}}_{n}^{2}$.
+By Proposition 3.1, the statistic $\hat{\mathcal{M}}^{2}_{n}$ is an unbiased estimator of $\mathcal{M}^{2}$ only under spherical symmetry.
+When the exact spherical symmetry does not hold, the bias of the kernel estimation with respect to the radius $\|Y\|$ is of order $h^{2}$, and it is of order $1/\kappa$ with respect to the direction $Y/\|Y\|$.
+Therefore, smaller values of $h$ and larger values of $\kappa$ are preferred to diminish the bias of the estimator $\hat{\mathcal{M}}_{n}^{2}$.
 We now turn to the weak convergence of $\hat{\mathcal{M}}_{n}^{2}$.
-Throughout this paper the symbol $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes weak convergence (convergence in distribution) of real valued random variables.
+Throughout this paper, the symbol $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes weak convergence (convergence in distribution) of real-valued random variables.
 Theorem 3.1.
-If 3.1 and 3.2 are satisfied, $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$, and the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
+If 3.1 and 3.2 are satisfied, $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$, and the distribution of $Y$ is spherically symmetric, that is $\mathcal{M}^{2}=0$, we have
 ${\hat{\mathcal{M}}^{2}_{n}\over s_{n}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,1),$
@@
 Note that the convergence rate in Theorem 3.1 is of order $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, which is asymptotically greater than $\sqrt{n}$ as $\kappa^{(p-1)/2}/(nh)\to 0$, by assumption.
-On the other hand, weak convergence with a rate $\sqrt{n}$ is obtained in the case where the distribution of $Y$ is not spherical symmetric.
+On the other hand, weak convergence with a rate $\sqrt{n}$ is obtained in the case where the distribution of $Y$ is not spherically symmetric.
 Theorem 3.2.
-Assume that 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$. If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
+Assume that 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$. If the distribution of $Y$ is not spherically symmetric, that is $\mathcal{M}^{2}>0$, we have
 $\sqrt{n}\big(\hat{\mathcal{M}}^{2}_{n}-\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]\big)\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2}),$
@@
 Remark 3.1.
-Theorem 3.2 can be used to derive the weak convergence
-$\sqrt{n}(\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2})$ (3.13)
-in the case, where the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$.
+Under exact sphericity (${\cal M}^{2}=0$), it follows from Theorem 3.1 that the variance of ${\cal M}^{2}_{n}$ is of order $\kappa^{(p-1)/2}/(n(n-1)h)$ and the estimator is consistent independently of the dimension.
+In the case ${\cal M}^{2}>0$, the situation is more complicated. Theorem 3.2 can be used to derive the weak convergence
+$\sqrt{n}(\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2}).$ (3.13)
 However, some care is necessary when replacing $\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]$ by $\mathcal{M}^{2}$ as the error of this replacement has to be of order $o(1/\sqrt{n})$.
@@
 However, a simple calculation shows that (3.14) and $\kappa^{(p-1)/2}/(nh)\to 0$ cannot simultaneously hold if $p\geq 4$.
-This issue can be mitigated via a simple bias reduction using the statistic
+This issue can be mitigated via a simple bias reduction
 $\tilde{\mathcal{M}}^{2}_{n}(a,\kappa)=\frac{1}{1-a}\hat{\mathcal{M}}^{2}_{n}(\kappa)-\frac{a}{1-a}\hat{\mathcal{M}}^{2}_{n}(a\kappa),$ (3.15)
-where $\hat{\mathcal{M}}^{2}_{n}(a\kappa)$ denotes the estimator $\hat{\mathcal{M}}^{2}_{n}$ using $a\kappa$ instead of $\kappa$ in the rapidly varying function $L(\cdot)$, see also Theorem 3 of Tsuruta, (2024).
-In addition, if we replace the kerne $K$ by its jackknife correction $\tilde{K}(x)=2\sqrt{2}K(\sqrt{2}x)-K(x)$, which satisfies $\phi_{2}(\tilde{K})=0$, we obtain from Proposition 3.1(ii)
+where $\hat{\mathcal{M}}^{2}_{n}(a\kappa)$ denotes the estimator $\hat{\mathcal{M}}^{2}_{n}$ using $a\kappa$ instead of $\kappa$ in the rapidly varying function $L(\cdot)$ in (2.10), see also Theorem 3 of TSURUTA2024105338.
+In addition, if we replace the kernel $K(\cdot)$ by its Jackknife correction $\tilde{K}(x)=2\sqrt{2}K(\sqrt{2}x)-K(x)$, which satisfies $\phi_{2}(\tilde{K})=0$, we obtain from Proposition 3.1(ii)
 $\mathbb{E}(\hat{\mathcal{M}}^{2}_{n})$ $={\mathcal{M}}^{2}+O(\kappa^{-3/2}+h^{3}),$ (3.16)
-and (3.13) holds under the conditions $\sqrt{n}/\kappa^{3/2}+\sqrt{n}h^{3}\to 0$ and $\kappa^{(p-1)/2}/(nh)\to 0$, which can be satisfied, whenever $p\leq 6$.
-If $p>6$ corresponding results hold under additional smoothness assumptions using the higher order expansions
+and (3.13) holds under the conditions $\sqrt{n}/\kappa^{3/2}+\sqrt{n}h^{3}\to 0$ and $\kappa^{(p-1)/2}/(nh)\to 0$, which can be satisfied, whenever $p\leq 6$, for example $\kappa=n^{\frac{3}{p+3}},h=n^{-\frac{3}{2(p+3)}}$.
+If $p>6$, corresponding results can be derived under additional smoothness assumptions using the higher order expansions
 $f(u,y)$ $=f(u,x)+\sum_{s=1}^{q}\frac{\theta^{s}}{s!}\xi^{\top}\mathcal{D}^{s}_{f}(u,x)\xi^{\otimes(s-1)}+O(\theta^{q+1}),$ (3.17)
 $f(u_{0}+h,v)$ $=f(u_{0},v)+\sum_{s=1}^{q}\frac{h^{s}}{s!}\left.\frac{\partial^{s}f(u,v)}{\partial u^{s}}\right|_{u=u_{0}}+O(h^{q+1}),$ (3.18)
-where $u^{\otimes s}$ denotes its $s$-th Kroneckerian power of the vector $u$.
-For example, if $q=3$, it follows observing $\int_{\Omega_{x}}\xi^{\top}\xi^{\otimes 2}\omega_{p-2}(\xi)=0$, and $\int x^{3}K(x)dx=0$ ($K$ is symmetry by Assumption 3.1) that the rate in (3.16) can be improved to $O(\kappa^{-2}+h^{4})$.
+where $u^{\otimes s}$ denotes the $s$-th Kroneckerian power of the vector $u$.
+For example, if $q=3$, it follows, observing $\int_{\Omega_{x}}\xi^{\top}\xi^{\otimes 2}\omega_{p-2}(\xi)=0$ and $\int x^{3}K(x)dx=0$ ($K$ is symmetric by Assumption 3.1), that the rate in (3.16) can be improved to $O(\kappa^{-2}+h^{4})$, and (3.13) holds under the conditions
+$\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}\to 0,$ (3.19)
+which can be satisfied, whenever $p\leq 8$, for example $\kappa=n^{\frac{3}{p+4}},h=n^{-\frac{3}{2(p+4)}}$.
 3.2 Some statistical applications
-By Proposition 3.1, Theorem 3.1 and 3.2, the statistic $\hat{\mathcal{M}}_{n}^{2}$ is a consistent estimator of the measure $\mathcal{M}^{2}$, which defines the deviation from sphericity.
+By Proposition 3.1, Theorem 3.1 and 3.2 and Remark 3.1, the statistic $\hat{\mathcal{M}}_{n}^{2}$ is a consistent estimator of the measure $\mathcal{M}^{2}$, which defines the deviation from sphericity.
 To use these results for uncertainty quantification, one requires estimates of the variances $s_{n}^{2}$ and $\sigma^{2}$ defined in (3.10) and (3.12), respectively.
@@
 $\hat{s}_{n}^{2}=\frac{\phi_{2}(K)\kappa^{(p-1)/2}}{2^{p-2}\pi^{(p-1)/2}n^{2}(n-1)^{2}c_{1}(\kappa)h^{2}}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j}),$ (3.20)
-which is obtained by replacing integral of the squared density in (3.10) by its corresponding estimate (2.8). We will use this estimate to define a test for exact sphericity (see Remark 3.2 below).
-The estimation of $\sigma^{2}$ is more difficult and we propose a jackknife approach for this purpose (see Chapter 5 of Lee,, 2003)).
+which is obtained by replacing the integral of the squared density in (3.10) by its corresponding estimate (2.8).
+We will use this estimate to define a test for exact sphericity (see Remark 3.2 below).
+The estimation of $\sigma^{2}$ is more difficult and we propose a Jackknife approach for this purpose (see Chapter 5 of Lee2003)).
 To be precise, let $\hat{\mathcal{M}}^{2}_{n-1}(-i)$ denote the estimator (2.12) of the minimum distance ${\cal M}^{2}$ based on the observations $Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}$, and define the pseudovalues
@@
 $\bar{\mathcal{M}}_{n}^{2}=n^{-1}\sum_{i=1}^{n}\tilde{\mathcal{M}}_{i}^{2}$ (3.23)
-denotes the jackknife estimate of the mean of $\mathcal{M}_{n}^{2}$.
+denotes the Jackknife estimate of the mean of $\tilde{\mathcal{M}}_{i}^{2}$.
 From Theorem 3.2 and Remark 3.1, we obtain a simple asymptotic confidence interval for $\mathcal{M}^{2}$, that is
@@
 Corollary 3.1.
-If the assumptions in Theorem 3.2 and in Remark 3.1 are satisfied, we have
+If the assumptions in Theorem 3.2 and (3.14) in Remark 3.1 are satisfied, we have
 $\lim_{n\to\infty}\mathbb{P}\big(\mathcal{M}^{2}\in\hat{I}_{n}\big)=1-\alpha.$
@@
 where $\Delta>0$ is a prespecified threshold.
-Note that this perspective of hypothesis testing aligns with the view expressed by Berger and Delampady, (1987), who argue that it is rare, and perhaps impossible, to have a null hypothesis that can be exactly modeled by a parameter being precisely 0.
-Similarly, Tukey, (1991), in the context of multiple comparisons of means, emphasizes that “All we know about the world teaches us that the effects of A and B are always different — in some decimal place — for any A and B.
+Note that this perspective of hypothesis testing aligns with the view expressed by berger1987, who argue that it is rare, and perhaps impossible, to have a null hypothesis that can be exactly modeled by a parameter being precisely 0.
+Similarly, tukey1991, in the context of multiple comparisons of means, emphasizes that “All we know about the world teaches us that the effects of A and B are always different — in some decimal place — for any A and B.
 Thus, asking ‘Are the effects different?’ is foolish.” We call hypotheses the form (3.25) and (3.26) relevant hypotheses in the following discussion.
-Note also that hypotheses of the form (3.25) have found considerable attention in equivalence testing in the field of biostatistics, which explains the notations $H_{0}^{\rm eq}$ and $H_{1}^{\rm eq}$ in (3.25) (see Wellek,, 2010) and that rejection of the null hypothesis in (3.25) allows to decide for approximate sphericity at a controlled type I error.
+Note also that hypotheses of the form (3.26) have found considerable attention in equivalence testing in the field of biostatistics, which explains the notations $H_{0}^{\rm eq}$ and $H_{1}^{\rm eq}$ in (3.26) (see wellek2010testing) and that rejection of the null hypothesis in (3.26) allows to decide for approximate sphericity at a controlled type I error.
 We propose to reject the null hypothesis in (3.25) if
@@
     For the test (3.27), we have
-    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}>\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}<\Delta.\end{cases}$ (3.28)
+    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}>\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}<\Delta.\end{cases}$ (3.29)
   • (b)
     For the test (3.28), we have
-    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}<\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}>\Delta.\end{cases}$ (3.29)
+    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}<\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}>\Delta.\end{cases}$ (3.30)
 We will illustrate the finite sample properties of the tests (3.27) and (3.28) in Section 5.
@@
   • (a)
-    An important question in testing hypotheses of the form (3.24) and (3.25) is the choice of the threshold $\Delta$, a similar argument can be given for the hypotheses (3.24).
-Note that this can be done in a data adaptive way.
-To be precise we concentrate on the hypotheses (3.25), and note these hypotheses are nested, that is if $H_{0}^{\rm eq}$ is satisfied for $\Delta_{1}$, then it is also satisfied for all $\Delta_{2}\leq\Delta_{1}$.
-The decision rules for different $\Delta$’s are nested in the same (if $H_{0}^{\rm eq}$ is rejected for $\Delta_{2}$, it also rejected for all $\Delta_{1}\geq\Delta_{2}$).
+    An important question in testing hypotheses of the form (3.25) and (3.26) is the choice of the threshold $\Delta$.
+In the following, we shall explain how to choose it in a data adaptive way for the hypotheses (3.26), and a similar argument can be given for the hypotheses (3.25).
+Note that these hypotheses are nested, that is, if $H_{0}^{\rm eq}$ is satisfied for $\Delta_{1}$, then it is also satisfied for all $\Delta_{2}\leq\Delta_{1}$.
+The decision rules for different $\Delta$’s are nested in the same way(if $H_{0}^{\rm eq}$ is rejected for $\Delta_{2}$, it is also rejected for all $\Delta_{1}\geq\Delta_{2}$).
 By the sequential rejection principle, we may simultaneously test the hypotheses in (3.26) for different $\Delta\geq 0$ starting at $\Delta=0$ and increasing $\Delta$ to find the minimum value of $\Delta$, say
-    $\hat{\Delta}_{\alpha}:=\min\big\{\Delta\geq 0\,|\,\hat{\cal M}^{2}_{n}\leq\Delta+u_{\alpha}\hat{s}_{n}/\sqrt{n}\big\}$ (3.30)
+    $\hat{\Delta}_{\alpha}:=\min\big\{\Delta\geq 0\,|\,\hat{\cal M}^{2}_{n}\leq\Delta+u_{\alpha}\hat{\sigma}_{n}/\sqrt{n}\big\}$ (3.31)
     for which $H_{0}$ in (3.26) is rejected.
@@
     $H_{0}^{\rm exact}:\mathcal{M}^{2}=0~~~~\text{versus}~~~~\mathcal{H}_{1}^{\rm exact}:\mathcal{M}^{2}>0~,$ (3.32)
-    it is worthwhile to mention that the theory developed so far also provides a test for these hypotheses. More precisely, we propose to reject the null hypothesis in (3.31), whenever
+    it is worthwhile mentioning that the theory developed so far also provides a test for these hypotheses without extra restrictions on the dimension $p$.
+More precisely, we propose to reject the null hypothesis in (3.32), whenever
     $\hat{\mathcal{M}}_{n}^{2}>u_{1-\alpha}{\hat{s}_{n}\over\sqrt{n}},$ (3.33)
-    where $\hat{s}_{n}$ is defined in (3.19). It then follows from Theorem 3.1 and 3.2 that this decision rule defines a consistent and asymptotic level $\alpha$ test.
+    where $\hat{s}_{n}$ is defined in (3.20).
+Then, it follows from Theorem 3.1 and 3.2 that this decision rule defines a consistent and asymptotic level $\alpha$ test.
+In Section 5.3, we compare the finite sample performance of the test (3.33) for the hypotheses (3.32) with other existing methods.
 4 Pivotal inference
-The statistical methodology developed in Section 3.2 requires estimation of the variance $\hat{s}_{n}^{2}$.
+The statistical methodology developed in Section 3.2 requires the estimation of the variance $\hat{s}_{n}^{2}$.
 In this section, we develop a pivotal confidence interval for the deviation $\mathcal{M}^{2}$ from sphericity and pivotal tests for the hypotheses (3.25) and (3.26).
@@
 $S_{n}(t)=\frac{\lfloor nt\rfloor}{2\sqrt{n}}\big(\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}-\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]\big),$ (4.1)
-whose asymptotic properties are investigated in following theorem.
-Throughout this section the symbol $\Rightarrow$ denotes weak convergence in the space $\ell^{\infty}([0,1])$ of bounded functions on the interval $[0,1]$.
+whose asymptotic properties are investigated in the following theorem.
+Throughout this section, the symbol $\Rightarrow$ denotes weak convergence in the space $\ell^{\infty}([0,1])$ of bounded functions on the interval $[0,1]$.
 Theorem 4.1.
@@
 The bandwidth condition (3.14) can be relaxed as discussed in Remark 3.1.
-We will now use the result (4.3) to develop pivotal inference for the measure $\mathcal{M}^{2}$. To be precise, we define the statistic
+We will now use the result (4.3) to develop pivotal inference for the measure $\mathcal{M}^{2}$. In particular, we define the statistic
 $\hat{V}_{n}=\int^{1}_{0}\big|\hat{\mathcal{M}}^{2}_{\lfloor nt\rfloor}-\hat{\mathcal{M}}^{2}_{n}\big|\,t\,\mathrm{d}t,$ (4.4)
@@
 $\hat{I}_{n}^{\rm piv}=\Big[\hat{\mathcal{M}}_{n}^{2}-\hat{V}_{n}q_{1-\alpha/2},\hat{\mathcal{M}}_{n}^{2}+\hat{V}_{n}q_{1-\alpha/2}\Big],$ (4.6)
-Similarly, we propose as a test for the hypotheses (3.24) and (3.25) to reject the null hypothesis, whenever
+Similarly, we propose tests for the hypotheses (3.25) and (3.26) to reject the null hypothesis, whenever
 $\hat{\mathcal{M}}_{n}^{2}$ $>\Delta+q_{1-\alpha}{\hat{V}_{n}},$ (4.7)
@@
 5 Finite sample properties
-In this section we illustrate the finite sample properties of the developed methodology by means of a small simulation study. All results presented here are based on $1000$ simulation runs.
+In this section, we illustrate the finite sample properties of the developed methodology through a small simulation study.
+For the choice of kernel functions, we use the Fisher-von-Mises distribution function(2.10) for $L(\cdot)$ and Epanechnikov kernel for $K(\cdot)$, i.e., $K(x)=\frac{3}{4}(1-x^{2})$, for $|x|<1$, and $K(x)=0$ otherwise.
+Both of the functions satisfy the Assumption 3.1.
+In addition, we implement the bias correction procedure in Remark 3.1 with $a=0.9$.
+All results presented here are based on $1000$ simulation runs.
 According to Theorem 3.2, the term $\kappa^{(p-1)/2}/(nh)$ controls the variance for the non-leading term in the Gaussian approximation, while $\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}$ accounts for the bias term with the bias reduction procedure discussed in Remark 3.2.
 In addition, for the sequential convergence, by Theorem 4.1, the inflated rate $(\log n)^{2}\kappa^{(p-1)/2}/(nh)$ controls the variance of the process of the non-leading term.
-In order to approximate Gaussianity in finite samples better, we recommend imposing a stronger control over the non-leading term, that is, $n^{3/2}\kappa^{(p-1)/2}/(nh)\to 0$.
+In order to approximate Gaussianity in finite samples better, we recommend imposing a stronger control over the non-leading term, that is, $n^{3/2}\times\kappa^{(p-1)/2}/(nh)$.
 Therefore, we consider bandwidths $(h,\kappa)$ satisfying
@@
 where SE denotes the standard error.
-In the following discussion we consider two models.
+In the following discussion, we consider two models.
 Model 1: $Y$ has a $3$-dimensional Gaussian distribution with mean vector $\mu=(1,0,2)^{\top}$ and covariance matrix
 $\Sigma=0.25\times\begin{pmatrix}1&0.3&0\\ 0.3&1&0\\ 0&0&1\end{pmatrix}$. (5.3)
-The deviation $\mathcal{M}^{2}$ was calculated numerically as $\mathcal{M}^{2}\approx 0.95$ , using a using a larger sample size $2500$ and averaged over $200000$ times of repeated simulation.
+Since it is difficult to obtain the analytical solution of $\mathcal{M}^{2}$, we calculate the deviation $\mathcal{M}^{2}$ numerically.
+We obtain $\mathcal{M}^{2}\approx 0.95$, using a larger sample size $2500$ and averaged over $20000$ times of repeated simulation.
 For the selection of the smoothing parameters via (5.2), we use
@@
 The Jackknife confidence interval achieves more accurate simulated coverage rates when the sample size is small, while the pivotal confidence intervals are more conservative for smaller sample sizes.
-A potential explanation of this observation is that in this case the widths of the pivotal method are much larger than those of the Jackknife method.
+A potential explanation for this observation is that when the sample size is small, the widths of the pivotal method are much larger than those of the Jackknife method.
 We observe similar patterns for Model 2.
@@
 5.2 Relevant hypothesis
-In Table 3 and 4, we display the simulated rejection rates of Jackknife test (3.27) and the pivotal test for the hypotheses (3.25) for different values of $\Delta$ and a significance level of $5\%$.
+In Table 3 and 4, we display the simulated rejection rates of Jackknife test (3.28) and the pivotal test (4.8) for the hypotheses (3.26) for different values of $\Delta$ and a significance level of $5\%$.
 For both models the results reflect the asymptotic properties of the test described in Corollary 3.2.
@@
 Table 3: Simulated rejection rates (in %) of the Jackknife test (3.28) and pivotal test (4.8) for the relevant hypotheses (3.26) in Model 1.
-In the “interior of the null hypothesis”, where $\mathcal{M}^{2}>\Delta$, the rejection rates of both tests decreases to 0 as $\Delta$ decreases and the sample size increases.
+In the “interior of the null hypothesis”, where $\mathcal{M}^{2}>\Delta$, the rejection rates of both tests quickly decrease to 0 as $\Delta$ decreases and the sample size increases.
+Note that this is a desirable behavior for a test of the composite hypotheses (3.26) as one has to assure for a type I error that
+$\sup_{{\cal M}^{2}\geq\Delta}\mathbb{P}(\text{ ``rejection''})\leq\alpha.$
+Thus, the level of the test is calibrated at the “boundary” of null hypothesis ${\cal M}^{2}=\Delta$ to be close to $\alpha$.
 Under the alternative, where $\mathcal{M}^{2}<\Delta$, the rejection rates of both tests increase to $1$ as $\Delta$ increases and the sample size increases.
@@
 1000 0.0 0.0 0.2 0.2 4.9 4.8 82.8 71.0 98.7 92.6 100.0 99.3
-Table 4: Simulated rejection rates (in %) of Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in Model 2.
-Summarizing, for testing for relevant deviations from sphericity from independent identically distributed data, the Jackknife test exhibits some advantages compared to the pivotal test if the sample size is sufficiently large.
+Table 4: Simulated rejection rates (in %) of Jackknife test (3.28) and the pivotal test (4.8) for the relevant hypotheses (3.26) in Model 2.
+In summary, for testing for relevant deviations from sphericity from independent identically distributed data, the Jackknife test exhibits some advantages compared to the pivotal test if the sample size is sufficiently large.
 However, we emphasize that this observation can only be made for independent data.
-In the case of dependencies, the Jackknife estimator does not yield a valid testing procedure.
-On the other hand, a careful inspection of the proofs in the appendix shows that the asymptotic statements in Section 4 remain valid for stationary processes under appropriate mixing (Bradley,, 2007), physical dependence (Wu,, 2005) or $m$-approximability (Hörmann and Kokoszka,, 2010) conditions.
+In the presence of dependencies, the Jackknife estimator does not yield a valid testing procedure.
+On the other hand, a careful inspection of the proofs in the appendix shows that the asymptotic statements in Section 4 remain valid for stationary processes under appropriate mixing (Bradley.2007), physical dependence (Wu2005) or $m$-approximability (HrmannKok) conditions.
 Consequently, Corollary 4.1 remains valid as well, and the pivotal test (4.8) has asymptotic level $\alpha$ and is consistent for the hypotheses (3.26).
@@
 1000 0 0 0.9 0.4 8.4 5.4 86.0 66.8 100.0 99.5 100.0 100.0
-Table 5: Simulated rejection rates (in %) of the Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in the case of dependent data.
-5.3 Data analysis
-For illustrate the potential applications of our approach , we analyze the log-returns of the daily exchange rate (close price) of the Yen to the Dollar and the Pound to the Euro from January 2, 2009, to December 31, 2009, which has also been investigated in Einmahl and Gantner, (2012) using a test for independent data.
-We downloaded the data from Yahoo Finance via the R package “quantmod” and obtained 260 entries for the log-returns, as the transactions of the currency can be conducted on weekdays including holidays.
+Table 5: Simulated rejection rates (in %) of the Jackknife test (3.28) and the pivotal test (4.8) for the relevant hypotheses (3.26) in the case of dependent data.
+5.3 Comparison with other methods
+In this section, we compare the minimum distance approach with alternative methods for the detection of deviations from sphericity.
+As all other papers focus on testing for exact ellipticity or sphericity, we consider the hypotheses (3.32) and the test (3.33) in Remark 3.2.
+Note that under the null hypothesis ${\cal M}^{2}=0$, the estimator $\hat{\mathcal{M}}_{n}^{2}$ is unbiased (see Proposition 3.1(i)).
+Therefore, no bias correction as discussed in Remark 3.1 is required in this situation.
+For the selection of bandwidths, we balance the squared bias under the alternative hypothesis, i.e., $\kappa^{-2}+h^{4}$ and the variance under Theorem 3.1, i.e., $\kappa^{(p-1)/2}/[n(n-1)h]$, and for $p=3$ use
+$h=1.5(n(n-1))^{-1/7}~,~~\kappa=1.5(n(n-1)h)^{1/3}.$ (5.7)
+In the following, we compare our test with other tests for the exact hypothesis (3.32) proposed by SCHOTT2002, MPQ2002, HP2007, PG2008, babic2021 and tang2024nonparametric.
+We consider the null hypothesis, where $Y$ follows a $3$-dimensional standard normal distribution.
+For the alternative we changed the first component of $Y$ to a $(\chi_{2}^{2}-2)/2$-distribution.
+In Table 6, we investigate the simulated rejection rate under the null hypothesis.
+The new test (3.33) provides a good approximation of the nominal level $\alpha=10\%$, $5\%$.
+The alternative tests exhibit a similar behavior under the null hypothesis, except for the test of tang2024nonparametric, which turns out to be conservative.
+The results under the alternative are displayed in Table 7.
+Although the new test (3.33) does not require Gaussianity or the existence of $4$-th moments, the empirical rejection rates under the alternative are larger than those of the tests HP2007, MPQ2002, SCHOTT2002, which require such assumptions, and comparable to the optimal parametric and semiparametric tests proposed by PG2008 and babic2021.
+Moreover, the new test (3.33) for the hypotheses (3.32) is more powerful in this scenario than the test based on kernel embedding of probabilities proposed by tang2024nonparametric.
+(3.32) TL HP MPQ
+[…]
+200 8.6 5.0 1.0 3.7 1.7 0.3 10.2 4.4 0.5 10.9 5.5 0.5
+300 10.1 5.5 1.2 3.9 1.9 0.2 10.1 5.2 1.3 10.6 5.4 1.2
+400 9.2 4.6 1.4 4.1 1.5 0.3 8.0 4.0 0.5 10.0 5.9 1.1
+(3.32) PG Schott BGH
+[…]
+200 8.6 5.0 1.0 11.0 5.1 1.2 10.1 4.8 0.6 10.9 4.7 0.9
+300 10.1 5.5 1.2 10.2 5.8 1.5 9.9 4.3 0.8 11.3 5.2 1.2
+400 9.2 4.6 1.4 12.2 7.2 1.3 10.8 5.4 1.2 10.9 5.9 0.9
+Table 6: Simulated rejection rates (in %) of the test (3.32) and the tests of tang2024nonparametric (TL), HP2007 (HP), MPQ2002 (MPQ), PG2008 (PG), SCHOTT2002 (Schott) and babic2021 (BGH) for the null hypothesis of exact sphericity.
+The sample sizes are $200$, $300$ and $400$.
+The nominal levels are 1010%, 55% and 11%.
+(3.32) TL HP MPQ
+[…]
+200 99.5 98.1 91.3 98.5 92.7 45.0 94.6 89.9 71.8 93.6 88.4 73.0
+300 99.9 99.9 99.8 100.0 99.9 93.8 99.9 99.0 94.8 99.4 98.9 94.6
+400 100.0 100.0 100.0 99.9 99.9 99.7 100.0 99.9 99.7 99.9 99.9 99.4
+(3.32) PG Schott BGH
+[…]
+200 99.5 98.1 91.3 100.0 100.0 100.0 75.2 64.2 42.3 100.0 100.0 99.9
+300 99.9 99.90 99.80 100.0 100.0 100.0 91.4 82.9 63.2 100.0 100.0 100.0
+400 100.0 100.0 100.0 100.0 100.0 100.0 95.7 91.6 79.0 100.0 100.0 100.0
+Table 7: Simulated rejection rates (in %) of the test (3.32) and the tests of tang2024nonparametric (TL), HP2007 (HP), MPQ2002 (MPQ), PG2008 (PG), SCHOTT2002 (Schott) and babic2021 (BGH) for the hypotheses of exact sphericity under the alternative hypothesis.
+The sample sizes are $200$, $300$ and $400$.
+The nominal levels are 1010%, 55% and 11%.
+5.4 Real data example
+To illustrate the potential applications of our approach, we analyze the log-returns of the daily exchange rate (close price) of the Yen to the Dollar and the Pound to the Euro from January 2, 2009, to December 31, 2009, which has also been investigated in einmahl2012 using a test for independent data.
+We downloaded the data from Yahoo Finance via the R package “quantmod” and obtained 260 entries for the log-returns, as the transactions of the currency can be conducted on weekdays, including holidays.
 The mean of the bivariate series turns out to be $9.4\times 10^{-5}$ and $2.7\times 10^{-4}$.
 In Figure 1, we plot the absolute autocorrelations of different lags of both of the series.
-We found that there exists some evidence of serial dependence, for example, the lag 12 of the Yen to Dollar and the lag 6 of Pound to Euro.
+We found that there exists some evidence of serial dependence, for example, the lag 12 of the Yen to Dollar and the lag 6 of the Pound to Euro.
 Figure 1: Absolute autocorrelations of the log-returns of the daily exchange rate of the Yen to the Dollar and the Pound to the Euro from January 2nd, 2009, to December 31st, 2009.
@@
 Appendix A Preliminary technical results
-In this section we present and prove several preliminary results, which will be essential for the proofs of the statements in Section 3 and 4.
+In this section, we present and prove several preliminary results, which will be essential for the proofs of the statements in Section 3 and 4.
 Our first result provides some properties of the Fisher-von-Mises distribution.
@@
 Proof.
-By equation (9) in Marco Di Marzio and Taylor, (2014), we have
+By equation (9) in DiMarzio03042014, we have
 $b_{j}(\kappa)\sim\frac{2^{j/2}\Gamma((p+j-1)/2)}{\kappa^{j/2}\Gamma((p-1)/2)}$ (5.11)
 as $\kappa\to\infty$ and with the notation $a_{j}(p)=2^{j/2}\Gamma((p+j-1)/2)/\Gamma((p-1)/2)$ the result (5.8) follows.
-For a proof of (5.8) we note that the modified Bessel function in (2.11) satisfies $\mathcal{I}_{v}(\kappa)\sim e^{\kappa}/\sqrt{2\pi\kappa}$, as $\kappa\to\infty$ (see equation (9.6.18) on Page 376 of Abramowitz and Stegun, (1968)).
+For a proof of (5.9) we note that the modified Bessel function in (2.11) satisfies $\mathcal{I}_{v}(\kappa)\sim e^{\kappa}/\sqrt{2\pi\kappa}$, as $\kappa\to\infty$ (see equation (9.6.18) on Page 376 of abramowitz1968handbook).
 Therefore, recalling the definition of the Fisher-von-Mises distribution in (2.10), we obtain
@@
 $c_{j}(\kappa)=\omega_{p-2}\int_{0}^{\pi}L^{j}(\kappa\cos\theta)(\sin\theta)^{p-2}\mathrm{d}\theta$ (5.37)
-and note that (see the proof of Theorem 3.1 in Marco Di Marzio and Taylor, (2014))
+and note that (see the proof of Theorem 3.1 in DiMarzio03042014)
 $\int_{\Omega_{v_{1}}}\xi\omega_{p-2}(\mathrm{d}\xi)=\mathbf{0}_{p},\quad\int_{\Omega_{v_{1}}}\xi\xi^{\top}\omega_{p-2}(\mathrm{d}\xi)={\omega_{p-2}\over(p-1)}(\mathbf{I}_{p}-v_{1}v_{1}^{\top}),$ (5.38)
@@
 $v_{1}^{\top}D^{2}_{f}(u,v_{1})v_{1}=0$ (5.39)
-(see the proof of Theorem 3.1 in Marco Di Marzio and Taylor,, 2014), which yields
+(see the proof of Theorem 3.1 in DiMarzio03042014), which yields
 $\int_{\Omega_{v_{1}}}\xi^{\top}D_{f}(u_{2},v_{1})\omega_{p-2}(\mathrm{d}\xi)$ $=0,$ (5.40)
@@
 $=\omega_{p-1}^{-1}(h^{*}(u)-h(u))(h^{*}(u)-h^{*}(u))=0.$ (5.87)
-Then, if $hf_{0}$ is the density of a spherical symmetric distribution (for some density $h$ on $\mathbb{R}^{+}$), we obtain for the squared $L^{2}$-distance between $f$ and $hf_{0}$
+Then, if $hf_{0}$ is the density of a spherically symmetric distribution (for some density $h$ on $\mathbb{R}^{+}$), we obtain for the squared $L^{2}$-distance between $f$ and $hf_{0}$
 $\Psi(h)$ $:=\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}(f(u,v)-h(u)f_{0}(v))^{2}\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v)$ (5.88)
@@
 where we have used Lemma A.2 three times in the last equality and Lemma A.1 for last two approximations.
-By Lemma B.1, ${\cal M}_{n}^{2}$ is a sum of martingale difference and we can apply a central limit theorem for sums of martingale differences (Theorem 1 in Hall,, 1984) to prove the statement in Theorem 3.1.
+By Lemma B.1, ${\cal M}_{n}^{2}$ is a sum of martingale difference and we can apply a central limit theorem for sums of martingale differences (Theorem 1 in HALL19841) to prove the statement in Theorem 3.1.
 We have already shown that $H_{n}$ is symmetric, $\mathbb{E}\{H_{n}(Y_{1},Y_{2})|Y_{1}\}=0$ a.s., and by Proposition 3.1 it follows that $\mathbb{E}\{H_{n}^{2}(Y_{1},Y_{2})\}<\infty$ for each $n$.
@@
 where
-$R_{21}$ $=\Big\|\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{-1}c_{1}^{-1}(\kappa)K\Big(\frac{u_{2}-U_{1}}{h}\Big)L(\kappa v_{2}^{\top}V_{1})f(u_{2},v_{2})\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|,$ (5.201)
+$R_{21}$ $=\Big\|\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{-1}c_{1}^{-1}(\kappa)K\Big(\frac{u_{2}-U_{1}}{h}\Big)L(\kappa v_{2}^{\top}V_{1})f(u_{2},v_{2})\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{2})-f(U_{1},V_{1})\Big\|,$ (5.202)
 $R_{22}$ $=\Big\|\omega_{p-1}^{-1}\Big\{\int_{\mathbb{R}^{+}}h^{-1}K\Big(\frac{u_{2}-U_{1}}{h}\Big)f_{U}(u_{2})\mathrm{d}u_{2}-f_{U}(U_{1})\Big\}\Big\|~,$ (5.203)
@@
 A Taylor expansion gives for the first term
-$R_{21}$ $=\Big\|\int_{-1}^{1}\int_{\mathcal{\mathbb{}}{S}^{p-1}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa v_{2}^{\top}V_{1})f(U_{1}+hu,v_{2})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|$ (5.204)
+$R_{21}$ $=\Big\|\int_{-1}^{1}\int_{\mathcal{\mathbb{}}{S}^{p-1}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa v_{2}^{\top}V_{1})f(U_{1}+hu,v_{2})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{2})-f(U_{1},V_{1})\Big\|$ (5.205)
 $\leq\Big\|\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{v_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)$ (5.206)
-$~~~~~~~~~~~~~~~~~~\times f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-f_{U,V}(U_{1},V_{1})\Big\|$ (5.206)
+$~~~~~~~~~~~~~~~~~~\times f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-f(U_{1},V_{1})\Big\|$ (5.207)
 $\lesssim\Big\|\int_{-1}^{1}\int_{0}^{\pi}\int_{\Omega_{V_{1}}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa\theta)\big\{f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)$ (5.208)
@@
 $\mathbb{E}[g(Y_{1})]=\mathcal{M}^{2}~~~\text{ and }~~\mathrm{Var}(g(Y_{1}))=\sigma^{2}.$ (5.215)
-Therefore, by Donsker’s Theorem (see the discussion on page 225 - 226 in Van Der Vaart and Wellner,, 1996) it follows that
+Therefore, by Donsker’s Theorem (see the discussion on page 225 - 226 in van1996weak) it follows that
 $\big\{S_{n}^{\circ}(t)\big\}_{t\in[0,1]}\Rightarrow\big\{\sigma\mathbb{B}(t)\big\}_{t\in[0,1]}$ (5.216)